Emacs/W3 User’s Manual

Third Edition, Emacs/W3 Version 4.0 June 1998

William M. Perry wmperry@cs.indiana.edu Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993 - 1995 William M. Perry Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996 - 1998 Free Software Foundation Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

W3

Users can browse the World Wide Web from within Emacs by using Emacs/W3. All of the widely used (and even some not very widely used) url schemes are supported, and it is very easy to add new methods as the need arises.

Emacs/W3 provides some core functionality that can be readily re-used from any program in Emacs. Users and other package writers are encouraged to Web-enable their applications and daily work routines with the library.

Emacs/W3 is completely customizable, both from Emacs-Lisp and from stylesheets See Chapter 4 [Stylesheets], page 19. If there is any aspect of Emacs/W3 that cannot be modified to your satisfaction, please send mail to the w3-beta@xemacs.org mailing list with any suggestions. See Appendix A [Reporting Bugs], page 62.

This manual corresponds to Emacs/W3 v4.0pre.46

1 Getting Started

If installed correctly, starting Emacs/W3 is quite painless. Just type M-x w3 in a running Emacs session. This will retrieve the default page that has been configured — by default the documentation for Emacs/W3 at Indiana University. The default homepage is specified

by the w3-default-homepage variable.

If the default page is not retrieved correctly at startup, you will have to do some customization.

Once started, you can use the mouse and the menu or use the following key commands (for more commands and more detail, see Chapter 2 [Basic Usage], page 5).

move forward press the space bar,

move backwards press the backspace key,

move to the next HTML reference on the page press the TAB key,

move to the previous HTML reference on the page press the SHIFT and TAB keys at the same time. If this does not work (some text terminals cannot distinguish between TAB and SHIFT-TAB, pressing the ALT and TAB keys should also work.

follow a link put the cursor over it and press the RETURN key, or click the left mouse button on it,

fetch a url press the Control and o keys at the same time, type the url, and then press the RETURN key,

return to the last URL you were at press the l key,

quit W3 mode press the q key.

1.1 Downloading

Emacs/W3 will work with Emacs 19.29 and later and XEmacs 19.14 and later, but if you’re using a 19.x Emacs then you will need to get the latest custom and widget libraries.

Emacs Available from the GNU archive ftp://prep.ai.mit.edu/pub/gnu or one of it’s many mirrors.

XEmacs Available from the XEmacs archive http://www.xemacs.org/ or one of it’s many mirrors.

Emacs/W3

          http://www.cs.indiana.edu/elisp/w3/docs.html is the main distribution

point for Emacs/W3.

Emacspeak A speech synthesizer package for Emacs and XEmacs. More information is

available at http://www.cs.cornell.edu/Info/People/raman/emacspeak/

1.2 Building and Installing

Emacs/W3 uses GNU ‘configure’ (see [Top], page 1) to control installation. configure will attempt to find what version of Emacs you have and where it is installed. If it finds both Emacs and XEmacs, then XEmacs is used (but see below for how to change this). Apart from the usual options, the following options are accepted:

‘--with-xemacs’

Use XEmacs.

‘--with-emacs’

Use Emacs.

‘--with-lispdir=dir’

Put lisp files (*.el and *.elc) in dir. If this is not specified, and neither is ‘--with-package-dir’, then the lisp files go into ‘EMACS/site-lisp’.

‘--with-package-dir=dir’

If using XEmacs, install Emacs/W3 as a package in dir. Please note that this and the ‘--prefix’ argument are mutually exclusive.

‘--with-makeinfo=makeinfo’

Use makeinfo to build info files from Texinfo files. ‘configure’ will normally find makeinfo if it’s available, you should only need to specify this if it’s not called makeinfo or if it isn’t in a directory in your ‘PATH’.

‘--with-custom=dir’

Use the custom package in dir. ‘configure’ will attempt to find a suitable custom package, you should not need to specify this yourself if the custom package is in Emacs’s load-path.

‘--enable-site-install’

Install Emacs/W3 for the site, rather than just yourself. This only affects

whether ‘make dotemacs’ affects ‘~/.emacs’ or ‘site-lisp/default.el’.

These are the most useful of the normal ‘configure’ options.

‘--prefix=dir’

This is the top level directory and by default everything is installed somewhere below this. This is ‘/usr/local’ by default.

‘--infodir=dir’

Where to put the info files. This is ‘prefix/info’ by default.

‘--data-dir=dir’

Where to put date files (default stylesheets). This is ‘prefix/share’ by default unless ‘--with-package-dir=pack-dir’ was given in which case they go into ‘pack-dir/etc/w3’.

The directory that the byte-compiled lisp files will be installed into is controlled by

the ‘--prefix’, ‘--with-package-dir’ and ‘--with-lispdir’ options. The directory for

the info and data files is likewise controlled by the ‘--prefix’, ‘--with-package-dir’ and

‘--infodir’ or ‘--data-dir’ options.

Normally these are the only things that need to be installed, they can by compiled by ‘make all’, or ‘make w3’ and ‘make info’. ‘make install’ will install the lisp, info and data files. ‘make all’ in the ‘texi’ directory will create the info files and also dvi files. html and postscript can be generated by ‘make html’ and ‘make ps’ respectively.

1.3 Startup Files

Emacs/W3 needs a directory for each user to store options, history and the cache. w3configuration-directory controls this directory, which is ‘~/.w3’ by default.

1.3.1 Emacs/W3 profile

Emacs/W3 keeps a file called ‘profile’ in your configuration directory that sets many variables. Warning: this file will overide any options that you set in your ‘.emacs’. You must either edit ‘profile’ directly or use w3-menu-save-options to save your settings.

If you prefer, you can set w3-default-configuration-file to specify a different configuration file. This file does not need to be dedicated to Emacs/W3 because Emacs/W3 will delimit its part of the file so you can set this to ‘.emacs’ if you want. However, while Emacs/W3 will save it’s options to the correct part of the file, it will read (and execute) the entire file when starting.

1.3.2 Default stylesheets

Emacs/W3 will look for style-sheets in w3-configuration-directory as well as the sitewide directories. In particular it will look for ‘dark.css’ or ‘stylesheet-dark’ if you’re using a dark background and ‘light.css’ or ‘stylesheet-light’ if you’re using a light

background as well as ‘stylesheet’ and ‘default.css’. If w3-default-stylesheet is not

nil then the file that it names will be used as well. For more information, See Chapter 4 [Stylesheets], page 19.

1.3.3 History

Emacs/W3 keeps a file called ‘history’ in the configuration directory. This is a list of all the links you have visited. You can change the file where the history is stored by setting url-global-history-file to the name of the file you’d prefer. See Section 3.7 [Global History], page 14.

1.3.4 Hotlists

Hotlists (sometimes called bookmarks, but not to be confused with Emacs’s bookmarks) are a list of urls. Emacs/W3 supports mosaic’s hotlist format which associates an alias with each url — See Section 3.5 [Hotlist Handling], page 13. The w3-hotlist-file variable specifies the hotlist file. It defaults to ‘.mosaic-hotlist-default’.

2 Basic Usage

Emacs/W3 is similar to the Info package all Emacs users hold near and dear to their hearts (See section “Info” in The Info Manual, for a description of Info). Basically, space and backspace control scrolling, and return or the middle mouse button follows a hypertext link. The tab and Meta-tab keys maneuver around the various links on the page.

NOTE: Starting with Emacs/W3 4.0, form entry areas in a page can be typed directly into. This is one of the main differences in navigation from version 2.0. If you are used to using the f and b keys to navigate around a buffer, I suggest training yourself to always use tab and M-tab — it will save time and frustration on pages with lots of form fields.

By default, hypertext links are surrounded by ’[[’ and ’]]’ on non-graphic terminals (VT100, DOS window, etc.). On a graphics terminal, the links are in shown in different colors. For information on how to change this, See Chapter 4 [Stylesheets], page 19.

There are approximately 50 keys bound to special Emacs/W3 functions. The basic rule of thumb regarding keybindings in Emacs/W3 is that a lowercase key takes an action on the current document, and an uppercase key takes an action on the document pointed to by the hypertext link under the cursor.

There are several areas that the keybindings fall into: movement, information, action, and miscellaneous.

2.1 Movement

All the standard Emacs bindings for movement are still in effect, with a few additions for convenience.

space Scroll downward in the buffer. With prefix arg, scroll down that many screenfuls

          (w3-scroll-up).

M-space Goes to next document.

M-del Goes to previous document.

backspace, C-?

Scroll upward in the buffer. With prefix arg, scroll up that many screenfuls

          (scroll-down).

< Goes to the start of document (w3-start-of-document).

> Goes to the end of document (w3-end-of-document).

Meta-tab, Shift-tab, b

Attempts to move backward one link area in the current document (w3-widget-backward). Signals an error if no previous links are found.

tab, f, n Attempts to move forward one link area in the current document (w3-widget-forward). Signals an error if no more links are found.

B, HB Takes one step back along the path in the current history (w3-history-backward). Has no effect if at the beginning of the session history.

F, HF Takes one step forward along the path in the current history (w3-history-forward). Has no effect if at the end of the session history.

l Return to the last buffer shown before this buffer (w3-goto-last-buffer).

q Kill this buffer (w3-quit).

Q, u Bury this buffer, but don’t kill it (w3-leave-buffer).

2.2 Information

These functions relate information about one or more links on the current document.

v This shows the url of the current document in the minibuffer (url-view-url).

V This shows the url of the hypertext link under point in the minibuffer (w3-

          view-this-url).

i Shows miscellaneous information about the currently displayed document (w3document-information). This includes the url, the last modified date, mime headers, the http response code, and any relationships to other documents. Any security information is also displayed.

I Shows information about the url at point (w3-popup-info).

s This shows the html source of the current document in a separate buffer (w3source-document). The buffer’s name is based on the document’s url.

S Shows the html source of the hypertext link under point in a separate buffer (w3-source-document-at-point). The buffer’s name is based on the document’s url.

k This stores the current document’s url in the kill ring, and also in the current window-system’s clipboard, if possible (w3-save-url).

K Stores the url of the document under point in the kill ring, and also in the current window-system’s clipboard, if possible (w3-save-this-url).

2.3 Action

First, here are the keys and functions that bring up a new hypertext page, usually creating a new buffer.

m Choose a link from the current buffer and follow it (w3-complete-link). A completing-read is done on all the links, so space and TAB can be used for completion.

return Pressing return when over a hyperlink attempts to follow the link under the

          cursor (w3-follow-link).

Pressing return when over a form input field can cause auto-submission of the form. This is for Mosaic and Netscape compatibility. If there is only one item in the form other than submit or reset buttons, then the form will be submitted.

Middle Mouse Button

Attempt to follow a hypertext link under the mouse cursor (w3-follow-mouse). Clicking on a form input field will prompt in the minibuffer for the data to insert into the input field. Type checking is done, and the data is only entered into the form when data of the correct type is entered (ie: cannot enter 44 for ’date’ field, etc).

Control Middle Mouse Button, Meta return

Tries to retrieve the inlined image that is under point (w3-follow-inlined-image). It ignores any form entry areas or hyperlinks, and blindly follows any inlined image. Useful for seeing images that are meant to be used as hyperlinks when not on a terminal capable of displaying graphics.

D  Download the url at point (w3-download-url-at-point).

d Download the current url (w3-download-this-url).

m Selects a destination from a list of all the hyperlinks in the current buffer (w3complete-link). Use space and tab to complete on the links.

r, g Reloads the current document (w3-reload-document). The position within the buffer remains the same (unless the document has changed since it was last retrieved, in which case it should be relatively close). This causes an unconditional reload from the remote server — the locally cached copy is not consulted.

R Redraws the buffer without reloading document (w3-refresh-buffer).

C-o Prompts for a url in the minibuffer, and attempts to fetch it (w3-fetch). If there are any errors, or Emacs/W3 cannot understand the type of link requested, the errors are displayed in a hypertext buffer.

o Opens a local file, interactively (w3-open-local). This prompts for a local file name to open. The file must exist, and may be a directory. If the requested file is a directory and url-use-hypertext-dired is nil, then a dired-mode buffer is displayed. If nonnil, then Emacs/W3 automatically generates a hypertext listing of the directory. The hypertext mode is the default, so that all the keys and functions remain the same.

M-s Save a document to the local disk as HTML Source, Formatted Text, LaTeX Source, or Binary (w3-save-as).

Hv, C-c C-b

Show the current session’s history list (w3-show-history-list). This takes all the links that are in that internal list, and formats them as hypertext links in a list, Section 3.7 [Global History], page 14

And here are the commands to move around between Emacs/W3 buffers:

l Goes to the last WWW buffer seen (w3-goto-last-buffer).

p Prints the current document (w3-print-this-url). Choose from several different formats to print: formatted text, html source, PostScript (with ps-print), or by using LaTeX and dvips). Section 11.2 [Printing], page 53.

P Prints out the url under point in a variety of formats (w3-print-url-under-point). Section 11.2 [Printing], page 53.

q Quits WWW mode (w3-quit). This kills the current buffer and goes to the most recently visited buffer.

u, Q This (w3-leave-buffer) is similar to w3-quit, but the buffer is not killed, it is moved to the bottom of the buffer list (so it is the least likely to show up as the default with switch-to-buffer). This is different from w3-goto-last-buffer in that it does not return to the last WWW page visited — it is the same as using switch-to-buffer — the buffer left in the window is fairly random.

2.4 Miscellaneous

? Shows help for Emacs/W3 (w3-help).

C-c C-v Show what version of Emacs/W3 you’re running (w3-version).

M-m Mails the current document to someone (w3-mail-current-document). Choose from several different formats to mail: formatted text, html source, PostScript, or LaTeX source. When the html source is mailed, then an appropriate <base> tag is inserted at the beginning of the document so that relative links may be followed correctly by whoever receives the mail.

M-M Mails the document pointed to by the hypertext link under point to someone (w3-mail-document-under-point). Choose from several different formats to mail: formatted text, html source, PostScript, or LaTeX source. When the html source is mailed, then an appropriate <base> tag is inserted at the beginning of the document so that relative links may be followed correctly by whoever receives the mail.

M-t, A-t gdj1: bound to url-list-processes. What does this do?

c Send a mail to the author of the current document (w3-mail-document-

author).
M-x w3-insert-formatted-url

Insert a fully formatted html link into another buffer (w3-insert-formatted-url). This gets the name and url of either the current buffer, or, with a prefix arg, of the link under point, and construct the appropriate <a...>...</a> markup and insert it into the desired buffer.

M-tab Inserts the url of the current document into another buffer (w3-insert-this-url'). Buffer is prompted for in the minibuffer. With prefix arg, uses the url of the link under point.

U Selects one of the <LINK> tags from this document and fetch it (w3-use-links). Links are attributes of a specific document, and can tell such things as who made the document, where a table of contents is located, etc. Link tags specify relationships between documents in two ways. Normal (forward) relationships (where the link has a REL="xxx" attribute), and reverse relationships (where the link has a REV="xxx" attribute). This first asks what type of link to follow (Normal or Reverse), then does a completing-read on only the links that have that type of relationship.

3 Compatibility with other Browsers

Due to the popularity of several other browsers, Emacs/W3 offers an easy transition to its much better way of life. This ranges from being able to share the same preferences files and disk cache to actually emulating the keybindings used in other browsers.

3.1 Emulation

Emacs/W3 can emulate the keybindings of lynx and netscape, but only one at a time. If you want emulation permanantly turned on, then you should add turn-on-lynx-emulation

or turn-on-netscape-emulation to w3-mode-hook.

3.2 Lynx emulation

turn-on-lynx-emulation turns on lynx emulation and turns off netscape emulation. lynx emulation is handled by the w3-lynx-emulation-minor-mode minor mode. For more information about lynx style hotlists, See Section 3.5 [Hotlist Handling], page 13.

:: work :: Document lynx emulation

Down arrow

Highlight next topic

Up arrow Highlight previous topic

Right arrow, Return, Enter

Jump to highlighted topic

Left arrow

Return to previous topic. gdj1: actually, this doesn’t seem to work quite right.

+ Scroll down to next page (Page-Down)

- Scroll up to previous page (Page-Up) SPACE Scroll down to next page (Page-Down) b Scroll up to previous page (Page-Up) C-a Go to first page of the current document (Home) C-e Go to last page of the current document (End) C-b Scroll up to previous page (Page-Up) C-f Scroll down to next page (Page-Down) C-n Go forward two lines in the current document C-p Go back two lines in the current document ) Go forward half a page in the current document (ignored) ( Go back half a page in the current document (ignored)

# Go to Toolbar or Banner in the current document, only works in XEmacs. gdj1:

is this what is meant by toolbar?

?, h Help (this screen) a Add the current link to a bookmark file c Send a comment to the document owner d Download the current link e Edit the current file (ignored) g Goto a user specified url or file i Show an index of documents (ignored) j Execute a jump operation (using hotlist) k Show a list of key mappings l List references (links) in current document m Return to main screen n Go to the next search string o Set your options p Print the current document q Quit r Delete hotlist entry s Enter a search string for an external search gdj1: really? u Go backwards in history / Search for a string within the current document v View a bookmark file V Go to the Visited Links Page x Force submission of form or link with no-cache z Cancel transfer in progress

[backspace]

Go to the history Page gdj1: really?

= Show file and link info \ Toggle document source/rendered view ! Spawn your default shell

* Toggle image links mode on and off (ignored) [ Toggle pseudo inlines mode on and off (ignorged) ] Send an http head request for the current doc or link (ignored)

C-r Reload current file and refresh the screen C-w Refresh the screen C-u Erase input line (ignored) C-g Cancel input or transfer C-t Toggle trace mode on and off (ignored) C-k Invoke the Cookie Jar Page (ignored)

3.3 Netscape emulation

turn-on-netscape-emulation turns on netscape emulation and turns off lynx emulation. netscape emulation is handled by the w3-netscape-emulation-minor-mode minor mode. For more information about netscape style hotlists, See Section 3.5 [Hotlist Handling], page 13.

M-a Add the current link to a bookmark file M-b Show hotlist file. M-f Search forward in document M-g Reapeat last search M-h Show history window M-i Load images M-l Goto a user specified url or file M-m Send mail M-n Open new frame M-o Open a local file M-p Print current document M-q Quit current document M-r Reload document and redraw M-s Save current document M-[left] Go back in history list

M-[right]

Go forward in history list

[left] Scroll left

[right] Scroll right

[up] Scroll up

[down] Scroll down

3.4 Masquerading

http allows servers to ask browsers what browser they are, and what version they are. Emacs/W3 allows you to choose the reply. There are functions to masquerade as lynx, netscape, ie or arena. For each browser there are three functions,

turn-on-browser-masquerade-mode,  turn-off-browser-masquerade-mode  and

w3-browser-masquerade-mode. The purpose of the first two is clear, w3-browser-masquerade-mode takes an optional argument which toggles the mode if it’s nil, turns off the mode if it’s 0 and turns the mode on otherwise.

If you’d prefer to masquerade as another browser, then you should call w3-masquerade-stub with three arguments: arg, app and version. arg has the same function as for w3browser-masquerade-mode, app is the name of the browser to masquerade as and version is the version.

Why would you want to masquerade as another browser when you could be advertising Emacs/W3? Well, some servers will only let certain browsers connect with them. This is cleary evil. Also some servers may alter what they present depending on the browser, this is probably a Good Thing but they might not know about Emacs/W3. Also one could argue that demanding the USER AGENT field is a breach of privacy, Emacs/W3 doesn’t have to send it (see Chapter 7 [Security], page 49), but the server doesn’t have to let you connect either.

3.5 Hotlist Handling

Emacs/W3 supports two types of hotlist, mosaic hotlists and html as used by lynx and netscape (which both call hotlists bookmarks). Unfortunately, not all hotlist operations are supported for html files at the moment.

In order to avoid having to traverse many documents to get to the same document over and over, Emacs/W3 supports a “hotlist” like Mosaic. This is a file that contains urls and aliases. Hotlists allow quick access to any document in the Web, providing it has been visited and added to the hotlist. The variable w3-hotlist-file determines where this information is saved. The structure of the file is compatible with Mosaic’s hotlist file, so

this defaults to ‘~/.mosaic-hotlist-default’.

Hotlist commands are:

ha Shows the hotlist entries matching a regular expression.

hi Adds the current document to the hotlist, with the buffer name as its identifier. Modifies the file specified by w3-hotlist-file. If this is given a prefixargument (via C-u), the title is prompted for instead of automatically defaulting to the document title.

hd Prompts for the alias of the entry to kill. Pressing the spacebar or tab will list out partial completions. The internal representation of the hotlist and the file specified by w3-hotlist-file are updated.

hr Some hotlist item names can be very unwieldy (‘Mosaic for X level 2 fill out form support’), or uninformative (‘Index of /’). Prompts for the item to rename in the minibuffer—use the spacebar or tab key for completion. After having chosen an item to rename, prompts for a new title until a unique title is entered. Modifies the file specified by w3-hotlist-file.

hu Prompts for the alias to jump to. Pressing the^{⟨}\underline{\overline{spacebar}}^{⟩} or^{⟨}\underline{\overline{tab}}^{⟩} key shows partial completions.

hv Converts the hotlist into html and displays it.

hA Appends another hotlist file to the one currently in memory.

hI Add the document pointed to by the hyperlink under point to the hotlist.

hR This rereads the default hostlist file specified by w3-hotlist-file.

3.6 History

Almost all web browsers keep track of the urls followed from a page, so that it can provide forward and back buttons to keep a path of urls that can be traversed easily.

If url-keep-history is non-nil, then Emacs/W3 keeps track of all the urls visited in an Emacs session. If t then Emacs/W3 will save the history list at the end of each session to the url-global-history-file file. The history list is simply a list of all the urls visited in the session.

To view a listing of the history for this session of Emacs/W3, use M-x w3-show-history-list (Hv) from any buffer, and Emacs/W3 generates an html document showing every url visited since Emacs started (or cleared the history list), and then format it. Any of the links can be chosen and followed to the original document. To clear the history list, choose ’Clear History’ from the ’Options’ menu.

Another twist on the history list mechanism is the fact that all Emacs/W3 buffers remember what url, buffer, and buffer position of the last document, and also keeps track of the next location jumped to from that buffer. This means that the user can go forwards and backwards very easily along the path taken to reach a particular document. To go forward, use the function w3-forward-in-history (F), to go backward, use the function

w3-backward-in-history (B). To view the entire history, use w3-show-history-list

(Hv).

3.7 Global History

Most web browsers also support the idea of a “history” of urls the user has visited, and it displays them in a different style than normal urls. Emacs/W3 will read and write history files generated by Emacs/W3, Mosaic v1 and v2 or netscape. Emacs/W3 looks at the file contents to determine the type of history.

If the variable url-keep-history is t, then Emacs/W3 keeps a list of all the urls visited in a session. The file is automatically written to disk every url-global-history-save-interval seconds and when exiting emacs. The list is added to those already in the file specified by url-global-history-file, which defaults to ‘~/mosaic.hst’ for MS operating systems, ‘~/mosaic.global-history’ for VMS and ‘~/.w3/history’ for everything else.

If any url in the list is found in the file, it is not saved, but new ones are added at the end of the file.

The function that saves the global history list is smart enough to notice what style of history list is being used (Netscape, Emacs/W3, or XMosaic), and writes out the new additions appropriately.

One of the nice things about keeping a global history files is that Emacs/W3 can use it as a completion table. When doing M-x w3-fetch, pressing the tab or space key will show all completions for a partial url. This is very useful, especially for very long urls that are not in a hotlist, or for seeing all the pages from a particular web site before choosing which to retrieve.

3.8 Display Variables

Emacs/W3 has many variable for you to fiddle with to get the display just right.

w3-display-frames

You can control what Emacs/W3 does with frame by setting w3-display-frames. It can be

nil Emacs/W3 will pretend not to understand frames at all.

as-nil Emacs/W3 will show hyperlinks to frames but will not fetch them (the same behaviour as lynx).

ask This is similar to as-nil, but Emacs/W3 will ask if you want to retrieve the frames.

t Emacs/W3 will dispaly the hyperlinks and fetch the frames.

w3-bullets

Emacs/W3 lets you decide what characters to use for bullets in unordered lists by setting w3-bullets. It is a association list, mapping list types to characters.

By default it is ((disc . ?*) (circle . ?o) (square . ?#) (none . ? )).
w3-echo-link

You can decide what should be displayed when tabbing through links by setting the w3-echo-link variable. It is a list and may contain the following symbols,

url Display the url of the target.

text Display the text of the link.

title Display the title attribute of the link.

name Display the name or id attribute of the link.

The default is (title url text name).
w3-horizontal-rule-char

Many html pages use horizontal lines (rules) to seperate sections of the page. You can control what character Emacs/W3 will use to draw these by setting w3-horizontal-rule-char. If it is a character (not a string) then Emacs/W3 will replicate that character across the screen, if it is nil then Emacs/W3 will use a terminal graphics character if possible. It is nil by default.

w3-use-terminal-characters

When Emacs/W3 draws table and rules, it needs to approximate line somehow.

If w3-use-terminal-characters it non-nil (the default) then Emacs/W3 will

use terminal graphics characters if they are available.

w3-use-terminal-characters-on-tty

Using terminal graphics characters on ttys will trigger display bugs in both XEmacs and FSF Emacs, but the display is usually readable with FSF Emacs.

w3-use-terminal-characters-on-tty controls whether to use terminal

graphics characters on ttys, it is nil by default.

w3-use-terminal-glyphs

Emacs/W3 can use glyphs rather than text properties for terminal graphics characters. Glyphs do not work with the most recent versions of XEmacs. This is t by default.

w3-defined-link-types

w3-defined-link-types is a list of names that have special significance as the values of ‘REL’ or ‘REV’ attributes of <link> elements. All members should be in lowercase.

w3-auto-image-alt

Some people do not feel it’s worth their time to add alt tags to their images, but Emacs/W3 can create alt tags on the fly for images that do not have them. To control this you can set w3-auto-image-alt to one of the following:

nil Do not create alt tags

string The string will be run through format with the filename of the image and so may have a single %s, for example "[IMAGE(%s)]"

function The function will be called with the filename of the images as the argument. This is the default, with w3-default-image-alt-func being the function.

w3-min-img-size, w3-default-image-alt-func, w3-dummy-img-alt-repl
w3-default-image-alt-func returns w3-dummy-img-alt-repl (‘*’ by

default) if the image’s height and width are both less than w3-min-img-size pixels (15 by default) and if the filename matches the w3-dummy-img-re regular expression. Otherwise, w3-default-image-alt-func returns the filename enclosed in a ‘[]’ pair.

w3-icon-format

Emacs/W3 will expect the standard icons to be in the format specified by w3-icon-format. This is gif by default, but could be xpm, xbm or any other format for that matter. It is added as a file extension to the icon name, but the variable’s value must be a symbol. If nil, then the server decides.

w3-delay-image-loads

You can choose whether Emacs/W3 retrieves images with the document, or delays loading them by setting w3-delay-image-loads. By default this is t if you compiled XEmacs with support for gifs, jpegs, pngs or imagick and nil otherwise.

w3-image-mappings

w3-image-mappings controls the mapping of mime types to image types for the ‘image’ package. Each entry is a cons cell of a mime type string and an image-type symbol.

w3-max-menu-length

Emacs/W3 will split menus into smaller submenus if they are longer than w3-

max-menu-length.
w3-max-menu-width

The maximum width of a pulldown menu choice.

w3-right-margin

Emacs/W3’s right margin is controlled by w3-right-margin. This is subtracted from (window-width) for each Emacs/W3 buffer and used as the fill-column. It is 2 by default.

w3-maximum-line-length

The maximum length of a line. If nil (the default) then lines can extend to the window margin.

w3-modeline-format

You can specify the modeline to use in ‘w3-mode’ by setting this.

w3-honor-stylesheets

If this is non-nil (the default) then Emacs/W3 will let a document specify a css stylesheet.

w3-user-colors-take-precedence

Emacs/W3 will ignore a document’s attempts to define certain colours if w3-

user-colors-take-precedence it non-nil. The default is nil.
w3-user-fonts-take-precedence

Emacs/W3 will ignore attempts by stylesheets or font tags to change certain fonts if this is non-nil.

3.8.1 Asynchronous behaviour

url-be-asynchronous

If this is non-nil then document retrievals over http will be down in the background.

url-default-retrieval-proc

This controls what happens when an asynchronous retrievel completes. It is url-default-callback by default but can be any function taking one argument. The argument specifies the file that has been retrieved. If there is no buffer associated with the file, then url-default-callback just puts a message in the minibuffer saying that the retrieval is complete, otherwise the action depends on the buffer.

w3-do-incremental-display

Emacs/W3 can de incremental display of pages if w3-do-incremental-display is t. It is nil by default.

w3-notify

You might want Emacs/W3 to notify you discreetly when it has finished preparing a page for your reading pleasure. You can control Emacs/W3’s behaviour in this situation by way of the w3-notify variable. It may take the following values:

newframe Puts the Emacs/W3 page in its own frame.

bully Make the Emacs/W3 page the current buffer and only window.

semibully

Make the Emacs/W3 page the current buffer in the same window. This is the default.

aggressive

Make the Emacs/W3 page the current buffer in the other window.

friendly Display the Emacs/W3 page in the other window, but don’t make it the current buffer.

polite Don’t display Emacs/W3 page, but print a message when ready and beep.

quiet The same as polite, but don’t beep.

meek Make no indication that the page is ready, in fact any other value is equivalent to meek.

4 Stylesheets

The way in which Emacs/W3 formats a document is very customizable. All formatting is now controlled by a default stylesheet set by the user with the w3-default-stylesheet variable. Emacs/W3 currently supports the W3C recommendation for Cascading Style Sheets, Level 1 (commonly known as CSS1) with a few experimental items from other W3C proposals. Wherever Emacs/W3 diverges from the specification, it will be clearly documented, and will be changed once a full standard is available.

Support for DSSSL is progressing, but spare time is at an all-time low. If anyone would like to help, please contact the author.

The following sections closely parallel the CSS1 specification so it should be very easy to look up what Emacs/W3 supports when browsing through the CSS1 specification. Please note that a lot of the text in the following sections comes directly from the specification as well.

4.1 Terminology

attribute HTML attribute, ie: ‘align=center’ — align is the attribute.

author The author of an HTML document.

block-level element An element which has a line break before and after (e.g. ’H1’ in HTML).

canvas The part of the UA’s drawing surface onto which documents are rendered.

child element A subelement in sgml terminology.

contextual selector A selector that matches elements based on their position in the document structure. A contextual selector consists of several simple selectors. E.g., the contextual selector ’H1.initial B’ consists of two simple selectors, ’H1.initial’ and ’B’.

css Cascading Style Sheets.

declaration A property (e.g. ’font-size’) and a corresponding value (e.g. ’12pt’).

designer The designer of a style sheet.

document html document.

element html element.

element type A generic identifier in sgml terminology.

fictional tag sequence A tool for describing the behavior of pseudo-classes and pseudo-elements. font size The size for which a font is designed. Typically, the size of a font is ap-

proximately equal to the distance from the bottom of the lowest letter with a descender to the top of the tallest letter with an ascender and (optionally) with a diacritical mark.

html extension Markup introduced by UA vendors, most often to support certain visual effects. The font, center and blink elements are examples of HTML extensions, as is the bgcolor attribute. One of the goals of css is to provide an alternative to html extensions.

inline element An element which does not have a line break before and after (e.g. ’strong’ in html)

intrinsic dimensions The width and height as defined by the element itself, not imposed by the surroundings. In this specification it is assumed that all replaced elements – and only replaced elements – come with intrinsic dimensions.

parent element The containing element in sgml terminology.

pseudo-element Pseudo-elements are used in css selectors to address typographical items (e.g. the first line of an element) rather than structural elements.

pseudo-class Pseudo-classes are used in css selectors to allow information external to the html source (e.g. the fact that an anchor has been visited or not) to classify elements.

property A stylistic parameter that can be influenced through css.

reader The person for whom the document is rendered.

replaced element An element that the css formatter only knows the intrinsic dimensions of. In html, img, input, textarea, select and object elements can be examples of replaced elements. E.g., the content of the img element is often replaced by the image that the src attribute points to. css1 does not define how the intrinsic dimensions are found.

rule A declaration (e.g. ’font-family: helvetica’) and its selector (e.g. ’H1’).

selector A string that identifies what elements the corresponding rule applies to. A selector can either be a simple selector (e.g. ’H1’) or a contextual selector (e.g. ’h1 b’) which consists of several simple selectors.

sgml Standard Generalized Markup Language, of which html is an application.

simple selector A selector that matches elements based on the element type and/or attributes, and not the element’s position in the document structure. E.g., ’H1.initial’ is a simple selector. style sheet A collection of rules.

ua User Agent, often a web browser or web client.

user Synonymous with reader.

weight The priority of a rule.

4.2 Basic Concepts

Designing simple style sheets is easy. One needs only to know a little HTML and some basic desktop publishing terminology. E.g., to set the text color of ’H1’ elements to blue, one can say:

       H1 { color: blue }

The example above is a simple CSS rule. A rule consists of two main parts: selector (’H1’) and declaration (’color: blue’). The declaration has two parts: property (’color’) and value (’blue’). While the example above tries to influence only one of the properties needed for rendering an HTML document, it qualifies as a style sheet on its own. Combined with other style sheets (one fundamental feature of CSS is that style sheets are combined) it will determine the final presentation of the document.

The selector is the link between the HTML document and the style sheet, and all HTML element types are possible selectors.

4.3 Pseudo-Classes/Elements

In css1, style is normally attached to an element based on its position in the document structure. This simple model is sufficient for a wide variety of styles, but doesn’t cover some common effects. The concept of pseudo-classes and pseudo-elements extend addressing in css1 to allow external information to influence the formatting process.

Pseudo-classes and pseudo-elements can be used in css selectors, but do not exist in the html source. Rather, they are "inserted" by the ua under certain conditions to be used for addressing in style sheets. They are referred to as "classes" and "elements" since this is a convenient way of describing their behavior. More specifically, their behavior is defined by a fictional tag sequence.

Pseudo-elements are used to address sub-parts of elements, while pseudo-classes allow style sheets to differentiate between different element types.

The only support pseudo-classes in Emacs/W3 are on the anchor tag (<a>...</a>).

User agents commonly display newly visited anchors differently from older ones. In css1, this is handled through pseudo-classes on the ’A’ element:

       A:link { color: red }       /* unvisited link */
       A:visited { color: blue }   /* visited links */
       A:active { color: lime }    /* active links */

All ’A’ elements with an ’HREF’ attribute will be put into one and only one of these groups (i.e. target anchors are not affected). UAs may choose to move an element from ’visited’ to ’link’ after a certain time. An ’active’ link is one that is currently being selected (e.g. by a mouse button press) by the reader.

The formatting of an anchor pseudo-class is as if the class had been inserted manually. A ua is not required to reformat a currently displayed document due to anchor pseudo-class transitions. E.g., a style sheet can legally specify that the ’font-size’ of an ’active’ link should be larger that a ’visited’ link, but the UA is not required to dynamically reformat the document when the reader selects the ’visited’ link.

Pseudo-class selectors do not match normal classes, and vice versa. The style rule in the example below will therefore not have any influence:

       A:link { color: red }
       <A CLASS=link NAME=target5> ... </A>

In css1, anchor pseudo-classes have no effect on elements other than ’A’. Therefore, the element type can be omitted from the selector:

       A:link { color: red }
       :link { color: red }

The two selectors above will select the same elements in CSS1.

Pseudo-class names are case-insensitive.

Pseudo-classes can be used in contextual selectors:

       A:link IMG { border: solid blue }

Also, pseudo-classes can be combined with normal classes:

       A.external:visited { color: blue }
       <A CLASS=external HREF="http://out.side/">external link</A>

If the link in the above example has been visited, it will be rendered in blue. Note that normal class names precede pseudo-classes in the selector.

4.4 The Cascade

In css, more than one style sheet can influence the presentation simultaneously. There are two main reasons for this feature: modularity and author/reader balance.

modularity A style sheet designer can combine several (partial) style sheets to reduce redundancy:

                 @import url(http://www.style.org/pastoral);
                 @import url(http://www.style.org/marine);
                 H1 { color: red }     /* override imported sheets */

author/reader balance Both readers and authors can influence the presentation through style sheets. To do so, they use the same style sheet language thus reflecting a fundamental feature of the web: everyone can become a publisher. The ua is free to choose the mechanism for referencing personal style sheets.

Sometimes conflicts will arise between the style sheets that influence the presentation. Conflict resolution is based on each style rule having a weight. By default, the weights of the reader’s rules are less than the weights of rules in the author’s documents. I.e., if there are conflicts between the style sheets of an incoming document and the reader’s personal sheets, the author’s rules will be used. Both reader and author rules override the ua’s default values.

The imported style sheets also cascade with each other, in the order they are imported, according to the cascading rules defined below. Any rules specified in the style sheet itself override rules in imported style sheets. That is, imported style sheets are lower in the cascading order than rules in the style sheet itself. Imported style sheets can themselves import and override other style sheets, recursively.

In css1, all ’@import’ statements must occur at the start of a style sheet, before any declarations. This makes it easy to see that rules in the style sheet itself override rules in the imported style sheets.

NOTE: The use of !important in css stylesheets is unsupported at this time.

Conflicting rules are intrinsic to the CSS mechanism. To find the value for an element/property combination, the following algorithm must be followed:

1. Find all declarations that apply to the element/property in question. Declarations apply if the selector matches the element in question. If no declarations apply, the inherited value is used. If there is no inherited value (this is the case for the ’HTML’ element and for properties that do not inherit), the initial value is used.

2. Sort the declarations by explicit weight: declarations marked ’!important’ carry more weight than unmarked (normal) declarations.

3. Sort by origin: the author’s style sheets override the reader’s style sheet which override the UA’s default values. An imported style sheet has the same origin as the style sheet from which it is imported.

4. Sort by specificity of selector: more specific selectors will override more general ones. To find the specificity, count the number of ID attributes in the selector (a), the number of CLASS attributes in the selector (b), and the number of tag names in the selector (c). Concatenating the three numbers (in a number system with a large base) gives the specificity. Some examples:

           LI            {...}  /* a=0 b=0 c=1 -> specificity =   1 */
           UL LI         {...}  /* a=0 b=0 c=2 -> specificity =   2 */
           UL OL LI      {...}  /* a=0 b=0 c=3 -> specificity =   3 */
           LI.red        {...}  /* a=0 b=1 c=1 -> specificity =  11 */
           UL OL LI.red  {...}  /* a=0 b=1 c=3 -> specificity =  13 */
           #x34y         {...}  /* a=1 b=0 c=0 -> specificity = 100 */

Pseudo-elements and pseudo-classes are counted as normal elements and classes, respectively.

5. Sort by order specified: if two rules have the same weight, the latter specified wins. Rules in imported style sheets are considered to be before any rules in the style sheet itself.

The search for the property value can be terminated whenever one rule has a higher weight than the other rules that apply to the same element/property combination.

This strategy gives author’s style sheets considerably higher weight than those of the reader. It is therefore important that the reader has the ability to turn off the influence of a certain style sheet, e.g. through a pull-down menu.

A declaration in the ’STYLE’ attribute of an element has the same weight as a declaration with an ID-based selector that is specified at the end of the style sheet:

     <STYLE TYPE="text/css">
       #x97z { color: blue }
     </STYLE>
     <P ID=x97z STYLE="color: red">

In the above example, the color of the ’P’ element would be red. Although the specificity is the same for both declarations, the declaration in the ’STYLE’ attribute will override the one in the ’STYLE’ element because of cascading rule number 5.

The ua may choose to honor other stylistic html attributes, for example ’ALIGN’. If so, these attributes are translated to the corresponding css rules with specificity equal to 1. The rules are assumed to be at the start of the author style sheet and may be overridden by subsequent style sheet rules. In a transition phase, this policy will make it easier for stylistic attributes to coexist with style sheets.

4.5 Properties

In the text below, the allowed values for each property are listed with a syntax like the following:

           Value: N | NW | NE
           Value: [ <length> | thick | thin ]{1,4}
           Value: <uri>? <color> [ / <color> ]?
           Value: <uri> || <color>

The words between < and > give a type of value. The most common types are <length>, <percentage>, <url>, <number>and <color> these are described in the section on [[units]]. The more specialized types (e.g. <font-family>and <border-style>) are described under the property where they appear.

Other words are keywords that must appear literally, without quotes. The slash (/) and the comma (,) must also appear literally.

Several things juxtaposed mean that all of them must occur, in the given order. A bar (|) separates alternatives: one of them must occur. A double bar (A || B) means that either A or B or both must occur, in any order. Brackets ([]) are for grouping. Juxtaposition is stronger than the double bar, and the double bar is stronger than the bar. Thus "a b | c || d e" is equivalent to "[ a b ] | [ c || [ d e ]]".

Every type, keyword, or bracketed group may be followed by one of the following modifiers:

• An asterisk (*) indicates that the preceding type, word or group is repeated zero or more times.

• A plus (+) indicates that the preceding type, word or group is repeated one or more times.

• A question mark (?) indicates that the preceding type, word or group is optional.

• A pair of numbers in curly braces ({A,B}) indicates that the preceding type, word or group is repeated at least A and at most B times.

Other than the value the following information is also shown. Supported Values:

If this is present, it lists the parts of the specification that Emacs/W3 currently supports.

Unsupported Values:

If this is present, it represents the parts of the specifcation that Emacs/W3 does not support.

Initial:

The default value for the property, unless explicitly set in a stylesheet.

Applies to:  What type of elements this property can be attached to.
Inherited:  Yes or no
Percentage values:  What a percentage value applies to when given.

4.5.1 Font Properties

Setting font properties will be among the most common uses of style sheets. Unfortunately, there exists no well-defined and universally accepted taxonomy for classifying fonts, and terms that apply to one font family may not be appropriate for others. E.g. ’italic’ is commonly used to label slanted text, but slanted text may also be labeled as being Oblique, Slanted, Incline, Cursive or Kursiv. Therefore it is not a simple problem to map typical font selection properties to a specific font.

The properties defined by CSS1 are described in the following sections.

4.5.1.1 font-family

Supported Values:

[[<family-name> | <generic-family>],]* [<family-name> | <genericfamily>]

Initial:  User specific
Applies to:  all elements
Inherited:  yes
Percentage values: N/A

The value is a prioritized list of font family names and/or generic family names. Unlike most other CSS1 properties, values are separated by a comma to indicate that they are alternatives:

       BODY { font-family: gill, helvetica, sans-serif }

There are two types of list values:

<family-name> The name of a font family of choice. In the last example, "gill" and "helvetica" are font families.

<generic-family> In the example above, the last value is a generic family name. The following generic families are defined: • ’serif’ (e.g. Times)

• ’sans-serif’ (e.g. Helvetica)

• ’cursive’ (e.g. Zapf-Chancery)

• ’fantasy’ (e.g. Western)

• ’monospace’ (e.g. Courier)

Style sheet designers are encouraged to offer a generic font family as a last alternative. Font names containing whitespace should be quoted:

       BODY { font-family: "new century schoolbook", serif }
       <BODY STYLE="font-family: 'My own font', fantasy">

If quoting is omitted, any whitespace characters before and after the font name are ignored and any sequence of whitespace characters inside the font name is converted to a single space.

4.5.1.2 font-style

Supported Values: normal | italic | oblique
Initial:  normal
Applies to:  all elements
Inherited:  yes
Percentage values: N/A

The ’font-style’ property selects between normal (sometimes referred to as "roman" or "upright"), italic and oblique faces within a font family.

A value of ’normal’ selects a font that is classified as ’normal’ in the UA’s font database, while ’oblique’ selects a font that is labeled ’oblique’. A value of ’italic’ selects a font that is labeled ’italic’, or, if that is not available, one labeled ’oblique’.

The font that is labeled ’oblique’ in the UA’s font database may actually have been generated by electronically slanting a normal font.

Fonts with Oblique, Slanted or Incline in their names will typically be labeled ’oblique’ in the UA’s font database. Fonts with Italic, Cursive or Kursiv in their names will typically be labeled ’italic’.

       H1, H2, H3 { font-style: italic }
       H1 EM { font-style: normal }

In the example above, emphasized text within ’H1’ will appear in a normal face.

4.5.1.3 font-variant

Value:  normal | small-caps
Initial:  normal
Applies to:  all elements
Inherited:  yes
Percentage values: N/A

Another type of variation within a font family is the small-caps. In a small-caps font the lower case letters look similar to the uppercase ones, but in a smaller size and with slightly different proportions. The ’font-variant’ property selects that font.

A value of ’normal’ selects a font that is not a small-caps font, ’small-caps’ selects a small-caps font. It is acceptable (but not required) in CSS1 if the small-caps font is a created by taking a normal font and replacing the lower case letters by scaled uppercase characters. As a last resort, uppercase letters will be used as replacement for a small-caps font.

The following example results in an ’H3’ element in small-caps, with emphasized words in oblique small-caps:

       H3 { font-variant: small-caps }
       EM { font-style: oblique }

There may be other variants in the font family as well, such as fonts with old-style numerals, small-caps numerals, condensed or expanded letters, etc. CSS1 has no properties that select those.

4.5.1.4 font-weight

Supported Values: normal | bold | 100 | 200 | 300 | 400 | 500 | 600 | 700 | 800 | 900

Unsupported Values:  bolder | lighter
Initial:  normal
Applies to:  all elements
Inherited:  yes
Percentage values:  N/A

The ’font-weight’ property selects the weight of the font. The values ’100’ to ’900’ form an ordered sequence, where each number indicates a weight that is at least as dark as its predecessor. The keyword ’normal’ is synonymous with ’400’, and ’bold’ is synonymous with ’700’. Keywords other than ’normal’ and ’bold’ have been shown to be often confused with font names and a numerical scale was therefore chosen for the 9-value list.

       P { font-weight: normal }   /* 400 */
       H1 { font-weight: 700 }     /* bold */

The ’bolder’ and ’lighter’ values select font weights that are relative to the weight inherited from the parent:

       STRONG { font-weight: bolder }

There is no guarantee that there will be a darker face for each of the ’font-weight’ values; for example, some fonts may have only a normal and a bold face, others may have eight different face weights. There is no guarantee on how a UA will map font faces within a family to weight values. The only guarantee is that a face of a given value will be no less dark than the faces of lighter values.

4.5.1.5 font-size

Supported Values:  <absolute-size> | <length>
Unsupported Values:  <percentage> | <relative-size>
Initial:  medium
Applies to:  all elements
Inherited:  yes
Percentage values:  relative to parent element’s font size

<absolute-size> An <absolute-size> keyword is an index to a table of font sizes computed and kept by the UA. Possible values are:

• xx-small

• x-small

• small

• medium

• large

• x-large

• xx-large

On a computer screen a scaling factor of 1.5 is suggested between adjacent indexes; if the ’medium’ font is 10pt, the ’large’ font could be 15pt. Different media may need different scaling factors. Also, the UA should take the quality and availability of fonts into account when computing the table. The table may be different from one font family to another.

<relative-size> A <relative-size> keyword is interpreted relative to the table of font sizes and the font size of the parent element. Possible values are larger or smaller. For example, if the parent element has a font size of ’medium’, a value of ’larger’ will make the font size of the current element be ’large’. If the parent element’s size is not close to a table entry, the UA is free to interpolate between table entries or round off to the closest one. The UA may have to extrapolate table values if the numerical value goes beyond the keywords.

Length and percentage values should not take the font size table into account when calculating the font size of the element.

Negative values are not allowed. On all other properties, ’em’ and ’ex’ length values refer to the font size of the current

element. On the ’font-size’ property, these length units refer to the font size of the parent element.

Note that an application may reinterpret an explicit size, depending on the context. E.g., inside a VR scene a font may get a different size because of perspective distortion.

Examples:

P { font-size: 12pt; }
BLOCKQUOTE { font-size: larger }
EM { font-size: 150% }
EM { font-size: 1.5em }

If the suggested scaling factor of 1.5 is used, the last three declarations are identical.

4.5.1.6 font

Value:

[ <font-style> || <font-variant> || <font-weight> ]? <font-size> [ / <line-height> ]? <font-family>

Initial:  not defined for shorthand properties
Applies to:  all elements
Inherited:  yes
Percentage values: allowed on <font-size> and <line-height>

The ’font’ property is a shorthand property for setting ’font-style’ ’font-variant’ ’fontweight’ ’font-size’, ’line-height’ and ’font-family’ at the same place in the style sheet. The syntax of this property is based on a traditional typographical shorthand notation to set multiple properties related to fonts.

For a definition of allowed and initial values, see the previously defined properties. Properties for which no values are given are set to their initial value.

       P { font: 12pt/14pt sans-serif }
       P { font: 80% sans-serif }
       P { font: x-large/110% "new century schoolbook", serif }
       P { font: bold italic large Palatino, serif }
       P { font: normal small-caps 120%/120% fantasy }

In the second rule, the font size percentage value (’80%’) refers to the font size of the parent element. In the third rule, the line height percentage refers to the font size of the element itself.

In the first three rules above, the ’font-style’, ’font-variant’ and ’font-weight’ are not explicitly mentioned, which means they are all three set to their initial value (’normal’). The fourth rule sets the ’font-weight’ to ’bold’, the ’font-style’ to ’italic’ and implicitly sets ’font-variant’ to ’normal’.

The fifth rule sets the ’font-variant’ (’small-caps’), the ’font-size’ (120% of the parent’s font), the ’line-height’ (120% times the font size) and the ’font-family’ (’fantasy’). It follows that the keyword ’normal’ applies to the two remaining properties: ’font-style’ and ’fontweight’.

4.5.2 Colors and Backgrounds

These properties describe the color (often called foreground color) and background of an element (i.e. the surface onto which the content is rendered). One can set a background color and/or a background image. The position of the image, if/how it is repeated, and whether it is fixed or scrolled relative to the canvas can also be set.

The ’color’ property inherits normally. The background properties do not inherit, but the parent element’s background will shine through by default because of the initial ’transparent’ value on ’background-color’.

NOTE: Currently, Emacs/W3 can only show background images under XEmacs. Emacs 19 doesn’t have the support in its display code yet.

4.5.2.1 color

Value:  <color>
Initial:  User specific
Applies to:  all elements
Inherited:  yes
Percentage values: N/A

This property describes the text color of an element (often referred to as the foreground color). There are different ways to specify red:

       EM { color: red }              /* natural language */
       EM { color: rgb(255,0,0) }     /* RGB range 0-255   */

See Section 4.6.3 [Color Units], page 42 for a description of possible color values.

4.5.2.2 background-color

Value:  <color> | transparent
Initial:  transparent
Applies to:  all elements
Inherited:  no
Percentage values: N/A

This property sets the background color of an element.

       H1 { background-color: #F00 }

4.5.2.3 background-image

Value:  <url> | none
Initial:  none
Applies to:  all elements
Inherited:  no
Percentage values: N/A

This property sets the background image of an element. When setting a background image, one should also set a background color that will be used when the image is unavailable. When the image is available, it is overlaid on top of the background color.

       BODY { background-image: url(marble.png) }
       P { background-image: none }

4.5.2.4 background-repeat

This property is not supported at all under Emacs/W3.

4.5.2.5 background-attachment

This property is not supported at all under Emacs/W3.

4.5.2.6 background-position

This property is not supported at all under Emacs/W3.

4.5.2.7 background

Value:

<background-color> || <background-image> || <background-repeat> || <background-attachment> || <background-position>

Initial:  not defined for shorthand properties
Applies to:  all elements
Inherited:  no
Percentage values: allowed on <background-position>

The ’background’ property is a shorthand property for setting the individual background properties (i.e., ’background-color’, ’background-image’, ’background-repeat’, ’backgroundattachment’ and ’background-position’) at the same place in the style sheet.

Possible values on the ’background’ properties are the set of all possible values on the individual properties.

       BODY { background: red }
       P { background: url(chess.png) gray 50% repeat fixed }

The ’background’ property always sets all the individual background properties. In the first rule of the above example, only a value for ’background-color’ has been given and the other individual properties are set to their initial value. In the second rule, all individual properties have been specified.

4.5.3 Text Properties

4.5.3.1 word-spacing

Supported Values:  normal
Unsupported Values:  <length>
Initial:  normal
Applies to:  all elements
Inherited:  yes
Percentage values:  N/A

The length unit indicates an addition to the default space between words. Values can be negative, but there may be implementation-specific limits. The UA is free to select the exact spacing algorithm. The word spacing may also be influenced by justification (which is a value of the ’align’ property).

       H1 { word-spacing: 0.4em }

Here, the word-spacing between each word in ’H1’ elements would be increased by ’1em’. NOTE: Emacs/W3 cannot currently support this, due to limitations in Emacs. It may

be implemented in the future.

4.5.3.2 letter-spacing

Supported Values:  normal
Unsupported Values:  <length>
Initial:  normal
Applies to:  all elements
Inherited:  yes
Percentage values:  N/A

The length unit indicates an addition to the default space between characters. Values can be negative, but there may be implementation-specific limits. The UA is free to select the exact spacing algorithm. The letter spacing may also be influenced by justification (which is a value of the ’align’ property).

       BLOCKQUOTE { letter-spacing: 0.1em }

Here, the letter-spacing between each character in ’BLOCKQUOTE’ elements would be increased by ’0.1em’.

NOTE: Emacs/W3 cannot currently support this, due to limitations in Emacs. It may be implemented in the future.

4.5.3.3 text-decoration

Supported Values:  none | underline | line-through | blink
Unsupported Values:  overline
Initial:  none
Applies to:  all elements
Inherited:  no, but see clarification below
Percentage values:  N/A

This property describes decorations that are added to the text of an element. If the element has no text (e.g. the ’IMG’ element in HTML) or is an empty element (e.g. ’<EM></EM>’), this property has no effect. A value of ’blink’ causes the text to blink.

The color(s) required for the text decoration should be derived from the ’color’ property value.

This property is not inherited, but elements should match their parent. E.g., if an element is underlined, the line should span the child elements. The color of the underlining will remain the same even if descendant elements have different ’color’ values.

       A:link, A:visited, A:active { text-decoration: underline }

The example above would underline the text of all links (i.e., all ’A’ elements with a ’HREF’ attribute).

NOTE: The ’line-through’ property is only supported under XEmacs currently. A patch has been sent to the Emacs maintainers to add support for this, but it has not made it into the main distribution yet.

4.5.3.4 vertical-align

This is currently unsupported in Emacs/W3.

4.5.3.5 text-transform

Supported Values:  none
Unsupported Values:  capitalize | uppercase | lowercase
Initial:  none
Applies to:  all elements
Inherited:  yes
Percentage values:  N/A

’capitalize’ Uppercases the first character of each word.

’uppercase’ Uppercases all letters of the element.

’lowercase’ Lowercases all letters of the element.

’none’ Neutralizes inherited value.

The actual transformation in each case is human language dependent.

       H1 { text-transform: uppercase }

The example above would put ’H1’ elements in uppercase text.

NOTE: This capability was in the previous version of Emacs/W3, but has not been reimplemented in the new display code yet. Please feel free to send me patches.

4.5.3.6 text-align

Value:  left | right | center | justify
Initial:  User specific
Applies to:  block-level elements
Inherited:  yes
Percentage values: N/A

This property describes how text is aligned within the element. The actual justification algorithm used is UA and human language dependent.

Example:

       DIV.center { text-align: center }

Since ’text-align’ inherits, all block-level elements inside the ’DIV’ element with ’CLASS=center’ will be centered. Note that alignments are relative to the width of the element, not the canvas.

4.5.3.7 text-indent

Not currently implemented in Emacs/W3.

4.5.3.8 line-height

Not currently implemented in Emacs/W3.

4.5.4 Box Properties

4.5.5 Classification

These properties classify elements into categories more than they set specific visual parameters.

The list-style properties describe how list items (i.e. elements with a ’display’ value of ’list-item’) are formatted. The list-style properties can be set on any element, and it will inherit normally down the tree. However, they will only be have effect on elements with a ’display’ value of ’list-item’. In HTML this is typically the case for the ’LI’ element.

4.5.5.1 display

Value:  block | inline | list-item | none
Extensions:  line
Initial:  inline
Applies to:  all elements

Inherited:  no
Percentage values: N/A

This property describes how/if an element is displayed on the canvas (which may be on a printed page, a computer display etc.).

An element with a ’display’ value of ’block’ opens whitespace suitable for a paragraph break. Typically, elements like ’H1’ and ’P’ are of type ’block’. A value of ’list-item’ is similar to ’block’ except that a list-item marker is added. In HTML, ’LI’ will typically have this value.

An element with a ’display’ value of ’inline’ results in a new inline box on the same line as the previous content.

A value of ’none’ turns off the display of the element, including children elements and the surrounding box.

       P { display: block }
       EM { display: inline }
       LI { display: list-item }
       IMG { display: none }

The last rule turns off the display of images.

A value of ’line’ results in a single line break. Emacs/W3 needs this extension to be able to fully specify the behaviour of br and hr elements within a stylesheet.

NOTE: Emacs/W3 defaults to using ’inline’ for this property, which is a slight deviation from the specification.

4.5.5.2 white-space

Value:  normal | pre | nowrap
Initial:  normal
Applies to:  block-level elements
Inherited:  yes
Percentage values: N/A

This property declares how whitespace inside the element is handled: the ’normal’ way (where whitespace is collapsed), as ’pre’ (which behaves like the ’PRE’ element in HTML) or as ’nowrap’ (where wrapping is done only through BR elements):

       PRE { white-space: pre }
       P   { white-space: normal }

4.5.5.3 list-style-type

Value: disc | circle | square | decimal | lower-roman | upper-roman | loweralpha | upper-alpha | none

Initial:  disc
Applies to:  elements with ’display’ value ’list-item’
Inherited:  yes
Percentage values: N/A

This property is used to determine the appearance of the list-item marker if ’list-style-image’ is ’none’ or if the image pointed to by the URL cannot be displayed.

Fo example:

       OL { list-style-type: decimal }       /* 1 2 3 4 5 etc. */
       OL { list-style-type: lower-alpha }   /* a b c d e etc. */
       OL { list-style-type: lower-roman }   /* i ii iii iv v etc. */

4.5.5.4 list-style-image

Value:  <url> | none
Initial:  none
Applies to:  elements with ’display’ value ’list-item’
Inherited:  yes
Percentage values: N/A

This property sets the image that will be used as the list-item marker. When the image is available it will replace the marker set with the ’list-style-type’ marker.

NOTE: This is currently unimplemented in Emacs/W3.

       UL { list-style-image: url(http://png.com/ellipse.png) }

4.5.5.5 list-style-position

Supported Values:  outside
Unsupported Values:  inside
Initial:  outside
Applies to:  elements with ’display’ value ’list-item’
Inherited:  yes
Percentage values:  N/A

The value of ’list-style-position’ determines how the list-item marker is drawn with regard to the content. For a formatting example see section 4.1.3.

4.5.5.6 list-style

Value:  <keyword> || <position> || <url>
Initial:  not defined for shorthand properties
Applies to:  elements with ’display’ value ’list-item’
Inherited:  yes
Percentage values: N/A

The ’list-style’ property is a shorthand notation for setting the three properties ’liststyle-type’, ’list-style-image’ and ’list-style-position’ at the same place in the style sheet.

       UL { list-style: upper-roman inside }
       UL UL { list-style: circle outside }
       LI.square { list-style: square }

Setting ’list-style’ directly on ’LI’ elements can have unexpected results. Consider:

       <STYLE TYPE="text/css">
         OL.alpha LI  { list-style: lower-alpha }
         UL LI        { list-style: disc }
       </STYLE>
       <BODY>
         <OL CLASS=alpha>

           <LI>level 1
           <UL>
              <LI>level 2
           </UL>
         </OL>
       </BODY>

Since the specificity (as defined in the cascading order) is higher for the first rule in the style sheet in the example above, it will override the second rule on all ’LI’ elements and only ’lower-alpha’ list styles will be used. It is therefore recommended to set ’list-style’ only on the list type elements:

       OL.alpha  { list-style: lower-alpha }
       UL        { list-style: disc }

In the above example, inheritance will transfer the ’list-style’ values from ’OL’ and ’UL’ elements to ’LI’ elements.

A URL value can be combined with any other value:

       UL { list-style: url(http://png.com/ellipse.png) disc }

In the example above, the ’disc’ will be used when the image is unavailable.

4.5.6 Media Selection

To specify that a stylesheet declaration should only apply when using a certain media type (ie: different font families preferred when printing versus on-screen presentation), the declarations should be wrapped in the proposed @media directive.

The @media directive takes two arguments, the media type, and a block of style declarations.

       @media print {
         BODY { font-size: 10pt }
         H1 { font-size: 14pt }
       }

The ’@media’ construct also allows to put include style sheet rules for various media in the same style sheet:

       @media print {
         BODY { font-size: 10pt }
       }
       @media screen {
         BODY { font-size: 12pt }
       }

Currently, the following media types are defined.

Print Output for paged opaque material, and for documents viewed on screen in print preview mode.

Screen A continuous presentation for computer screens.

Projector Paged presentation for projected presentations.

Braille For braille tactile feedback devices.

Speech Aural presentation.

Light The stylesheet will only be applied if the user is using a light background.

Dark The stylesheet will only be applied if the user is using a dark background.

Emacs The stylesheet will only be applied if the user is running in Emacs 19.

XEmacs The stylesheet will only be applied if the user is running in XEmacs 19.

All The default value, the style sheet applies to all output devices.

4.5.7 Speech Properties

Those of us who are sighted are accustomed to visual presentation of html documents, frequently on a bitmapped display. This is not the only possible presentation method, however. Aural presentation, using a combination of speech synthesis and ’audio icons’, provides an alternative presentation. This form of presentation is in current use by the blind and print-impaired communities.

Often such aural presentation occurs by converting the document to plain text and feeding this to a ’screen reader’ – software or hardware that simply reads all the characters on the screen. This results in less effective presentation than would be the case if the document structure were retained.

There are other large markets for aural presentation, including in-car and home entertainment use; aurual or mixed aural/visual presentation is thus likely to increase in importance over the next few years. Realizing that that the aural rendering is essentially independent of the visual rendering:

• Allows orthogonal aural and visual views.

• Allows browsers to optionally implement both aural and visual views to produce truly multimodal documents.

4.5.7.1 volume

Value:  <percentage> | mute | x-soft | soft | medium | loud | x-loud
Initial:  medium
Applies to:  all elements
Inherited:  yes
Percentage values: relative to user-specified mapping

The legal range of percentage values is 0% to 100%. There is a fixed mapping between keyword values and percentages:

• ’x-soft’ = ’0%’

• ’soft’ = ’25%’

• ’medium’ = ’50%’

• ’loud’ = ’75%’

• ’x-loud’ = ’100%’

Volume refers to the median volume of the waveform. In other words, a highly inflected voice at a volume of 50 might peak well above that. Note that ’0%’ does not mean the same as "mute". 0% represents the minimum audible volume level and 100% corresponds to the maximum comfortable level. The UA should allow the values corresponding to 0% and 100% to be set by the user. Suitable values depend on the equipment in use (speakers, headphones), the environment (in car, home theater, library) and personal preferences. Some examples:

• A browser for in-car use has a setting for when there is lots of background noise . 0% would map to a fairly high level and 100% to a quite high level. The overall values are likely to be human adjustable for comfort, for example with a physical volume control: what this proposal does is adjust the dynamic range.

• Another speech browser is being used in the home, late at night, (don’t annoy the neighbors) or in a shared study room. 0% is set to very quiet and 100% to a fairly quiet level, too. As with the first example, there is a low slope; the dynamic range is reduced. The actual volumes are low here, wheras they were high in the first example.

• In a quiet and isolated house, an expensive hifi home theatre setup. 0% is set fairly low and 100% to quite high; there is wide dynamic range.

The same authors stylesheet could be used in all cases, simply by mapping the 0 and 100 points suitably at the client side.

4.5.7.2 pause-before

Value:  <time> | <percentage>
Initial:  UA specific
Applies to:  all elements
Inherited:  no
Percentage values: speed

This property specifies the pause before elements. It may be given in an absolute units (seconds, milliseconds) or as a relative value in which case it is relative to the reciprocal of the ’speed’ property: if speed is 120 words per minute (ie a word takes half a second – 500 milliseconds) then a pause-before of 100% means a pause of 500 ms and a pause-before of 20% means 100ms.

Using relative units gives more robust stylesheets in the face of large changes in speed.

4.5.7.3 pause-after

Value:  <time> | <percentage>
Applies to:  all elements
Inherited:  no
Percentage values: speed

This property specifies the pause after elements. Values are specified the same way as ’pause-before’.

4.5.7.4 pause

Value:  [<time> | <percentage> ]{1,2};
Applies to:  all elements
Inherited:  no
Percentage values: speed

The ’pause’ property is a shorthand for setting ’pause-before’ and ’pause-after’. The first value is pause-before and the second is pause-after. If only one value is given, it applies to both properties.

Examples:

       H1 { pause: 20ms }       /* pause-before: 20ms; pause-after: 20ms */
       H2 { pause: 30ms 40ms }  /* pause-before: 30ms; pause-after: 40ms */
       H3 { pause-after: 10ms } /* pause-before: ?;    pause-after: 10ms */

4.5.7.5 cue-before

Value:  <url> | none
Initial:  none
Applies to:  all elements
Inherited:  no

Auditory icons are another way to distinguish semantic elements. Sounds may be played before, and/or after the element to delimit it. The same sound can be used both before and after, using the cue property.

Examples:

       A  { cue-before: url(bell.aiff); cue-after: url(dong.wav) }
       H1 { cue-before: url(pop.au); cue-after: url(pop.au) }
       H1 { cue: url(pop.au) }  /* same as previous */

4.5.7.6 cue-after

See Section 4.5.7.5 [cue-before], page 39.

4.5.7.7 cue

See Section 4.5.7.5 [cue-before], page 39.

4.5.7.8 cue-during

Value:  <url> | mix | none
Initial:  mix
Applies to:  all elements
Inherited:  no

Similar to the cue-before and cue-after properties, this indicates sound to be played during an element as a background (ie the sound is mixed in with the speech).

Examples:

       BLOCKQUOTE.sad { cue-during: url(violins.aiff) }

4.5.7.9 speed

Value:

<words-per-minute> | x-slow | slow | medium | fast | x-fast | faster | slower

Initial:  medium

Applies to:  all elements
Inherited:  yes

Specifies the speaking rate. Note that both absolute and relative keyword values are allowed (compare with Section 4.5.1.4 [font-weight], page 27).

4.5.7.10 voice-family

Value:

[[<specific-voice> | <generic-voice>],]* [<specific-voice> | <genericvoice>]

Initial:  device-specific
Applies to:  all elements
Inherited:  yes

The value is a prioritized list of voice family names. Generic families are male, female, and child.

Examples of specific voice families are: comedian, paul, lisa

Examples

       H1 { voice-family: announcer, male }
       P.part.romeo { voice-family: romeo, male }
       P.part.juliet { voice-family: juliet, female }

4.5.7.11 pitch

4.5.7.12 pitch-range

4.5.7.13 stress

Value:  <percentage>
Initial:  medium
Applies to:  all elements
Inherited:  yes

Specifies the level of stress (assertiveness or emphasis) of the speaking voice. English is a stressed language, and different parts of a sentence are assigned primary, secondary or tertiary stress. The value of property ’stress’ controls the amount of inflection that results from these stress markers.

Increasing the value of this property results in the speech being more strongly inflected. It is in a sense dual to property ’pitch-range’ and is provided to allow developers to exploit higher-end auditory displays.

4.5.7.14 richness

Value:  <percentage>
Initial:  medium (50%)

Applies to:  all elements
Inherited:  yes

Specifies the richness (brightness) of the speaking voice. Different speech devices may require the setting of one or more device-specific parameters to achieve this effect.

The effect of increasing richness is to produce a voice that carries – reducing richness produces a soft, mellifluous voice.

4.5.7.15 speak-punctuation

Value:  code | none
Initial:  none
Applies to:  all elements
Inherited:  yes

’code’ indicates that punctuation such as semicolons, braces, and so on are to be spoken literally. The default value of ’none’ means that punctuation is not spoken but instead is rendered naturally as various pauses.

4.5.7.16 speak-date

Value:  myd | dmy | ymd | none
Initial:  none
Applies to:  all elements
Inherited:  no

This is a hint that the element contains a date and also how that date should be spoken. month-day-year is common in the USA, while day-month-year is common in Europe and year-month-day is also used.

This should really be an HTML tag not a stylesheet property, since it gives semantic information about the content.

4.5.7.17 speak-numeral

Value:  digits | continous
Initial:  none
Applies to:  all elements
Inherited:  yes

4.5.7.18 speak-time

Value:  24 | 12 | none
Initial:  none
Applies to:  all elements
Inherited:  yes

4.6 Units

4.6.1 Length Units 4.6.2 Percentage Units

4.6.3 color Units

4.6.4 URLs

4.6.5 Angle Units

These are the legal angle units:

• deg: degrees

• grad

• rad: radians

4.6.6 Time Units

These are the legal time units:

• ms: milliseconds

• s: seconds

5 Supported URLs

::WORK:: List supported URL types, specific RFCs, etc.

Uniform Resource Locators (urls) are a specific form of Uniform Resource Identifiers (uri) described in rfc2396 which updates rfc1738 and rfc1808.

rfc2016 defines uniform resource agents.

uri have the form scheme:scheme-specific-part, where scheme appears in the menu below for urls supported by Emacs/W3.

ftp, nfs, http, https, rlogin, telnet, tn3270, irc and gopher urls all have the form scheme://[userinfo@]hostname[:port][/path] where ‘[’ and ‘]’ delimit optional parts. userinfo sometimes takes the form userinfo:password but you should beware of the security risks of sending cleartext passwords. hostname may be a domain name or a dotted decimal address. If the ‘:port’ is omitted then Emacs/W3 will use the well known port for that service. With the possible exception of telnet, it is very rare for ports to be specified, and it is possible using a non-standard port may have undesired consequences if a different service is listening on that port (eg. a gopher url specifying the smtp port can cause mail to be sent), but See Section 11.6 [Other Variables], page 55. The meaning of the path component depends on the service.

5.1 File urls

This allows Emacs/W3 to read arbitary files from hosts. Compressed files are handled, but support is hard-coded so that jka-compr-compression-info-list and so on have no affect. Suffixes recognized are ‘.z’, ‘.gz’ and ‘.Z’. If the url points to a directory, then it will try to retrieve a file named by url-directory-index-file (‘index.html’ by default) and parse it, otherwise you get a directory listing in ‘dired’ mode. If the url refers to a file on a remote host, then Emacs/W3 uses ‘ange-ftp’ (see [ange-ftp], page 1) or ‘efs’ (see [efs], page 1) to retrive the file. ftp:// and file:// are synonymous for Emacs/W3.

5.2 ftp urls

For details of usage see Section 5.1 [file], page 43. In Emacs/W3 file and ftp urls are synonymous and files on the localhost are retrieved directly rather than by ftp. Emacs/W3 relies on ‘ange-ftp’ (see [ange-ftp], page 1) or ‘efs’ (see [efs], page 1) to do the actual transfers.

5.3 nfs urls

gdj1: have I misunderstood the point of this? Since nfs is fairly transparent to the user (at least when it’s working), there isn’t very much to say here. An nfs url is similar to a file url except that it points to a file on a remote host that is handled by the automounter on the local host. The variable url-nfs-automounter-directory-spec may need to be tweaked depending on local configuration. The nfs url is defined in rfc2224.

5.4 info

Info urls are not an officially recognised url (gdj1: is this right?), but Emacs/W3 will parse them to produce a reference to a TeXinfo node, or ‘Top’ if one is not specified. ‘Info-mode’ will be used to browse the document.

5.5 http/https

The HyperText Transfer Protocol is the protocol used to get documents from the World Wide Web. Emacs/W3 supports http version 1.0 as defined in rfc1945 — now superseded by version 1.1 defined in rfc2068.

   If url-honor-refresh-requests is nil then ‘Refresh’ headers will not be honoured,

if t then they will always be honoured, otherwise the user will be asked for each request. The default is t. url-be-anal-about-file-attributes controls whether http is used to discover file attributes, or whether they’re just guessed. The default is nil which means that Emacs/W3 will make educated guesses.

https is a secure version of http defined in rfc2069 (gdj1: ?). Emacs/W3 requires ssl to handle this, See Appendix D [Installing SSL], page 68.

5.6 mailto

A mailto url will send an email message to the address in the url, for example ‘mailto:foo@bar.com’ would compose a message to foo@bar.com. Emacs/W3 uses the command specified by the url-mail-command variable to compose the email, this is urlmail by default which uses Gnu’s ‘message’ mode (see [Top], page 1) if available, otherwise the standard mail command. An X-Url-From header field containing the url of the document that contained the mailto url is added, as is an X-Mailer header field containing the version of Emacs/W3 being used.

rfc2368 extends the definition of mailto urls in rfc1738. The form of a mailto url is ‘mailto:mailbox[?header=contents[&header=contents]]’ where an arbitary number of headers can be added. If the header is ‘body’, then contents is put in the body otherwise a header header field is created with contents as its contents. Note that Emacs/W3 does not consider any headers as ‘dangerous’ so you should check them before sending the message.

Email messages are defined in rfc822.

5.7 mailserver

A mailserver url allows you to send an email to a person, but this url optionally specifies a subject and a body. The basic format is ‘mailserver:[mailbox/subject[/body]’. Thus,

‘mailserver:foo@bar.com/wibble/flibble’ will compose a message to foo@bar.com with

subject as the subject and body already in the body of the email. Note that both the subject and the body are hexed, but the subject cannot contain newlines.

5.8 news/nntp/snews

If the url doesn’t specify a host, then the host in url-news-server will be used, and unless the url has a port the news port as defined in url-default-ports (119) will be used. The username and password specified in the url will be used if present. The url may contain a message-id, in which case that article is displayed; it may contain a newsgroup in which case Gnus is used to display the newsgroup; or it may by empty in which case Gnus is called with no arguments. Emacs/W3 requires Gnus v5.x or Red, Quassia or Pterodactyl Gnus, See section “Top” in Gnus. The variable url-news-use-article-mode controls the displaying of news articles; if non-nil then articles are displayed in Gnus article mode, otherwise they are turned into html and rendered by Emacs/W3.

An nntp url is the same as a news url, except that the url may specify an article by its number.

5.9 rlogin/telnet/tn3270

To handle rlogin, telnet and tn3270 urls, Emacs/W3 runs an rlogin, telnet or tn3270 session (the program names and arguments are hardcoded) in a terminal-emulator buffer. Well-known ports are used if the url does not specify a port.

5.10 irc

Internet Relay Chat (irc) is handled by handing off the irc session to a function named in url-irc-function. This function must take five argumenst, host, port, channel, user and password. The channel argument specifies the channel to join immediately, this can be nil. By default this is url-irc-zenirc which processes the arguments and lets zenirc handle the session.

5.11 data

Data urls contain mime data in the url itself, by default the data is 8bit encoded ‘text/plain’, but the url can specify either or both the content-type and the encoding. Emacs/W3 will parse the url’s data as mime and display it appropriately. See Chapter 6 [MIME Support], page 47.

5.12 gopher

Gopher (Go for) was in someways the precurser to the world wide web and is becoming rarer as the web becomes more powerful. Nevertheless, there are still many gopher sites around and Emacs/W3 supports this protocol (of course). The variable url-gopher-labels maps gopher types to something else (gdj: ?) for displaying the gopher menus. url-gopher-icons maps gopher types to pictures. url-gopher-to-mime maps gopher types to mime types. If url-use-hypertext-gopher is non-nil, then gopher pages will be turned into html for Emacs/W3 to parse and display normally, otherwise Emacs/W3 will let ‘gopher.el’ handle all gopher requests which will lose gopher+ support and inlined searching. This is t by default.

5.13 finger

Finger urls will finger a given user at a given host, or ‘localhost’ if no host is specified, processing the results to create an html page for Emacs/W3 to display.

5.14 netrek

This is unsupported at present.

6 MIME Support

mime is an emerging standard for multimedia mail. It offers a very flexible typing mechanism. The type of a file or message is specified in two parts, separated by a ’/’. The first part is the general category of the data (text, application, image, etc.). The second part is the specific type of data (postscript, png, jpeg, etc.). So ‘text/html’ specifies an html document, whereas ‘image/x-xwindowdump’ specifies an image of an Xwindow taken with the ‘xwd’ program.

This typing allows much more flexibility in naming files. http/1.0 servers can now send back content-type headers in response to a request, and not have the client second-guess it based on file extensions. html files can now be named ‘something.png’ (not a great idea, but possible).

6.1 Adding MIME types based on file extensions

For some protocols however, it is still necessary to guess the content of a file based on the file extension. This type of guess-work should only be needed when accessing files via ftp, local file access, or old http/0.9 servers.

Instead of specifying how to view things twice, once based on content-type and once based on the file extension, it is easier to map file extensions to MIME content-types. The variable that controls this is mm-mime-extensions.

This variable is an assoc list of file extensions and the corresponding MIME content-type. A sample entry looks like: ‘(".movie" . "video/x-sgi-movie")’ This makes all files that end in ‘.movie’ (‘foo.movie’ and ‘bar.movie’) be interpreted as SGI animation files. If a content-type is defined for the document, then this is over-ridden. Regular expressions can NOT be used.

Both Mosaic and the NCSA http daemon rely on a separate file for mapping file extensions to MIME types. Instead of having the users of Emacs/W3 duplicate this in lisp, this file can be parsed using the url-parse-mimetypes function. This function is called each time w3 is loaded. It tries to locate mimetype files in several places. If the environment variable MIMETYPES is nonempty, then this is assumed to specify a UNIX-like path of mimetype files (this is a colon separated string of pathnames). If the MIMETYPES environment variable is empty, then Emacs/W3 looks for these files:

 1. ‘~/.mime-types’
 2. ‘/etc/mime-types’
 3. ‘/usr/etc/mime-types’
 4. ‘/usr/local/etc/mime-types’
 5. ‘/usr/local/www/conf/mime-types’

Each line contains information for one http type. These types resemble MIME types. To add new ones, use subtypes beginning with x-, such as application/x-myprogram. Lines beginning with # are comment lines, and suitably ignored. Each line consists of:

type/subtype ext1 ext2 ... extn

type/subtype is the MIME-like type of the document. ext* is any number of spaceseparated filename extensions which correspond to the MIME type.

6.2 Specifying Viewers

Not all files look as they should when parsed as an html document (whitespace is stripped, paragraphs are reformatted, and lots of little changes that make the document look unrecognizable). Files may be passed to external programs or Emacs Lisp functions to be viewed.

Not all files can be viewed accurately from within an Emacs session (PNG files for example, or audio files). For this reason, the user can specify file "viewers" based on MIME content-types. This is done with a standard mailcap file. See Appendix E [Mailcap Files], page 69.

As an alternative, the function mm-add-mailcap-entry can also be used from an appropriate hook. See Section 11.5 [Hooks], page 55. This functions takes three arguments, the major type ("image"), the minor type ("png"), and an assoc list of information about the viewer. Please see the url documentation for more specific information on what this assoc list should look like.

7 Security

There are an increasing number of ways to authenticate a user to a web service. Emacs/W3 tries to support as many as possible. Emacs/W3 currently supports:

Basic Authentication The weakest authentication available, not recommended if serious security is necessary. This is simply a string that looks like ‘user:password’ that has been Base64 encoded, as defined in RFC 1421.

Digest Authentication Jeffery L. Hostetler, John Franks, Philip Hallam-Baker, Ari Luotonen, Eric W. Sink, and Lawrence C. Stewart have an internet draft for a new authentication mechanism. For the complete specification, please see draft-ietf-http-digest-aa-01.txt in the nearest internet drafts archive^{1}.

SSL Encryption SSL is the Secure Sockets Layer interface. Emacs/W3 supports http transfers over an SSL encrypted channel, if the appropriate files have been installed. See Appendix D [Installing SSL], page 68.

7.1 Privacy

Sometimes you don’t want people to know who you are, or where you’ve been. http is quite happy to tell everyone it meets who you are and where you’ve come from. urlprivacy-level can be used to set how much information is given, it can be a list of the following symbols

‘email’ Do not send email address. This just sets url-personal-mail-address to nil.

‘os’ Do not send operating system

‘lastloc’ Do not send the last location

‘agent’ Do not send the User-Agent string (for an alternative approach, see Section 3.4 [Masquerading], page 13).

‘cookie’ Never accept cookies (see Chapter 8 [Cookies], page 50)

Alternatively url-privacy-level can be a single symbol,

‘none’ Send all information.

‘low’ Don’t send the last location. Equivalent to (lastloc)

‘high’ Don’t send the email address or last location. Equivalent to (email lastloc)

‘paranoid’

Don’t send anything. Equivalent to (email os lastloc agent cookie)

   If you change url-privacy-level then you should also call url-setup-privacy-info

to make sure that the changes propogate.

1 One is ftp://ds.internic.net/internet-drafts

8 Cookies

http is a stateless protocol which means that the server sees every request for pages independently with no idea of how it relates to any other request. Therefore the server has no idea whether or not you’ve seen a page before, or whether you’ve registered (if that’s an option). Cookies^{1} are used to add state to http sessions. Cookies are defined in rfc2109.

Cookies are saved in the file specified in url-cookie-file, which is w3-configuration-directory/cookies by default. Note that this file should probably not be world writable, and possibly not even world readable.

Some people see cookies as an invasion of privacy while others see them as a product of badly designed websites and buggy servers. Emacs/W3 lets you unconditionally reject all cookies by adding cookie to url-privacy-level or setting it to paranoid (see Chapter 7 [Security], page 49) but for those who want finer control over what to accept and reject,

Emacs/W3 offers url-cookie-trusted-urls and url-cookie-untrusted-urls which are

lists of regular expressions that match urls from which cookies should be accepted and rejected respectively. If a url matches patterns in both of these, then Emacs/W3 decides whether to accept or not based on the most specific match (the most specific match being the shortest match). Note that Emacs/W3 only considers the first match for each variable, so the regular expressions should be in increasing order of generality.

For even more control over which cookies are accepted, you can set url-cookie-confirmation to non-nil, in which case every time a cookie is offered Emacs/W3 will ask if you want to accept it. This only applies to cookies that would otherwise be accepted, Emacs/W3 will still reject cookies from urls matched in url-cookie-untrusted-urls.

1 In computer terms a cookie is data that a program holds but which has no meaning in itself. Cookies are not processed by the program (indeed the program may not even know what data they hold or what format it’s in) but is passed to libraries or servers which do understand it.

9 Non-Unix Operating Systems

9.1 VMS

:: WORK :: VMS Specific instriuctions

9.2 OS/2

:: WORK :: OS/2 Specific instructions

9.3 MS-DOS

:: WORK :: DOS Specific instructions

9.4 Windows

:: WORK :: 32bit Windows Specific instructions Chapter 10: Speech Integration 52

10 Speech Integration

:: WORK :: Emacspeak integration

11 Advanced Features

11.1 Disk Caching

A cache stores the information on a page on the local machine. When requesting a page that is in the cache, Emacs/W3 can retrieve the page from the cache more quickly than retrieving the page again from its location out on the network. With a well-populated cache, browsing the web is dramatically faster.

The first time a page is requested, Emacs/W3 retrieves the page from the network. When requesting a page that is in the cache, Emacs/W3 checks to see if the page has changed since it was last retrieved from the remote machine. If it has not changed, the local copy is used, saving the transmission of the file over the network.

To turn on disk caching, set the variable url-automatic-caching to non-nil, or choose the ’Caching’ menu item (under ‘Options’). That is all there is to it. Running the cleancache shell script fist is recommended, to allow for future cleaning of the cache. This shell script will remove all files that have not been accessed since it was last run. To keep the cache pared down, it is recommended that this script be run from at or cron (see the manual pages for crontab(5) or at(1) for more information)

With a large cache of documents on the local disk, it can be very handy when traveling, or any other time the network connection is not active (a laptop with a dial-on-demand PPP connection, etc). Emacs/W3 can rely solely on its cache, and avoid checking to see if the page has changed on the remote server. In the case of a dial-on-demand PPP connection, this will keep the phone line free as long as possible, only bringing up the PPP connection when asking for a page that is not located in the cache. This is very useful for demonstrations as well. To turn this feature on, set the variable url-standalone-mode to non-nil, or choose the ‘Use Cache Only’ menu item (under ‘Options’)

url-cache-expired decides whether or not a cache entry has expired. It is a function that take two times as it parameters and returns non-nil if the second time is “too old” when compared with the first time. url-cache-ignored-protocols is a list of protocols that

will never be cached, this is '("www" "about" "https" "mailto") by default. url-cache-

directory sets the directory to store the cache files, ‘"w3-configuration-directorycache/"’ by default. url-cache-creation-function sets the type of cache to use, it is a function that takes a url as an argument and returns the absolute pathname of the cachefile corresponding to that url. You may write your own function or use one of the two

ready built functions, url-cache-create-filename-using-md5 and url-cache-create-
filename-human-readable. The advantage of url-cache-create-filename-using-md5

is that there are very few cache collisions but is only “suitably fast” if you’re not using

XEmacs. url-cache-create-filename-human-readable will give a filename more obvi-

ously connected to the url, but it is more likely to conflict with other files.

11.2 Printing

If you want to print an html document, then Emacs/W3 needs to convert it into something that can be printed. You can choose from html source

This will simply print the raw html source code using lpr-buffer. An appropriate <base> tag is inserted at the beginning of the document.

Formatted text This will print the rendered document using lpr-buffer; so the conversion is handled by Emacs. This will print plain ASCII.

Postscript This will call the function in w3-postscript-print-function, which is ps-print-buffer-with-faces by default. This just tells Emacs to generate postscript as best it can.

LaTeX Emacs/W3 can generate a LaTeX equivalent of the html document. w3-print-command contains a command string to print ‘dvi’ files. It is ‘lpr -h -d’ by default. There are several variables controlling what the final LaTeX document looks like. :: WORK :: Document the new LaTeX backend

w3-latex-use-latex2e

If non-nil, configures the LaTeX engine to use the LaTeX2e syntax. A nil value indicates that LaTeX 2.0.9 compabibility will be used instead.

w3-latex-docstyle

The document style to use when printing or mailing converted html files in LaTeX.

w3-latex-packages

List of LaTeX packages to include. Currently this is only used if

          w3-latex-use-latex2e is non-nil.
w3-latex-use-maketitle

If non-nil, the LaTeX engine will use real LaTeX title pages for document titles.

w3-latex-print-links

If non-nil, prints the urls of hypertext links as endnotes at the end of the document. If set to footnote, prints the url’s as footnotes on each page.

11.3 Interfacing to Mail/News

More and more people are including urls in their signatures, and within the body of mail messages. It can get quite tedious to type these into the minibuffer to follow one.

With the latest versions of VM (the 5.9x series of betas) and Gnus (5.x), urls are automatically highlighted, and can be followed with the mouse or the return key. How the urls are viewed is determined by the variable browse-url-browser-function, and it should be set to the symbol browse-url-w3.

To access urls from within RMAIL, the following hook should do the trick.

     (add-hook 'rmail-mode-hook
       (function
        (lambda ()
          (define-key rmail-mode-map [mouse-2] 'w3-maybe-follow-link-mouse)
          (define-key rmail-mode-map "\r"      'w3-maybe-follow-link))))

11.4 Debugging HTML

For those people that are adventurous, or are just as anal as I am about people writing valid html, set the variable w3-debug-html to t and see what happens. Alternatively, you can set it to style to warn about stylistic issues as well. The debugging information will be written to the buffer named by w3-debug-buffer, ‘*HTML Debug*’ by default. To control font-lock highlighting in the html error buffer, use w3-html-errors-font-lock-keywords. After Emacs/W3 has displayed html errors for a page, it runs w3-display-errors-hook.

If a Emacs/W3 thinks it has encountered invalid html, then a debugging message is displayed.

:: WORK :: Need to list the different values w3-debug-html can have, and :: WORK :: what they do ::

gdj1: Does this refer to the macro? And if so, why?

11.5 Hooks

These are the various hooks that can be used to customize some of Emacs/W3’s behavior. They are arranged in the order in which they would happen when retrieving a document. These are all ’normal hooks’ in standard Emacs-terminology, meaning they are functions (or lists of functions) that are called consecutively.

w3-load-hook

These hooks are run the first time a url is fetched. All the Emacs/W3 variables are initialized before this hook is run.

w3-mode-hook

These hooks are run after a buffer has been parsed and displayed, but before any inlined images are downloaded and converted.

w3-source-file-hook

These hooks are run after displaying a document’s source.

11.6 Miscellaneous variables

There are lots of variables that control the real nitty-gritty of Emacs/W3 that the beginning user probably shouldn’t mess with. Here they are.

url-bad-port-list

List of ports to warn the user about connecting to. Defaults to just the mail, nntp and chargen ports so a malicious html author cannot spoof mail or news to other people.

url-confirmation-func

What function to use for asking yes or no functions. Possible values are 'yesor-no-p or 'y-or-n-p, or any function that takes a single argument (the prompt), and returns t only if a positive answer is gotten. Defaults to 'yes-

or-no-p.
url-passwd-entry-func

This is a symbol indicating which function to call to read in a password. If this variable is nil at startup, it is initialized depending on whether EFS or ange-ftp is being used. This function should accept the prompt string as its first argument, and the default value as its second argument.

url-max-password-attempts

When a protected document is requested, Emacs/W3 will prompt for a password. url-max-password-attempts controls how many attempts should be allowed, it is 5 by default.

w3-reuse-buffers

Determines what happens when w3-fetch is called on a document that has already been loaded into another buffer. Possible values are: nil, yes, and no. nil will ask the user if Emacs/W3 should reuse the buffer (this is the default value). A value of yes means assume the user wants to always reuse the buffer. A value of no means assume the user always wants to re-fetch the document.

url-show-status

Whether to show progress messages in the minibuffer. url-show-status controls if a running total of the number of bytes transferred is displayed. This Can cause a large performance hit if using a remote X display over a slow link, or a terminal with a slow modem.

mm-content-transfer-encodings

An assoc list of Content-Transfer-Encodings or Content-Encodings and the appropriate decoding algorithms for each. If the cdr of a node is a list, then this specifies the decoder is an external program, with the program as the first item in the list, and the rest of the list specifying arguments to be passed on the command line. If using an external decoder, it must accept its input from stdin and send its output to stdout. If the cdr of a node is a symbol whose function definition is non-nil, then that encoding can be handled internally. The function is called with 2 arguments, buffer positions bounding the region to be decoded. The function should completely replace that region with the unencoded information. Currently supported transfer encodings are: base64, x-gzip, 7bit, 8bit, binary, x-compress, x-hqx, and quoted-printable.

url-uncompressor-alist

An assoc list of file extensions and the appropriate uncompression programs for each. This is used to build the Accept-encoding header for http/1.0 requests.

w3-do-scripting

If this is non-nil then Emacs/W3 will do clien-side scripting. This is nil by default.

url-external-retrieval-program, url-external-retrieval-args

url-external-retrieval-program names the external program that is run to retrieve urls. It is ‘www’ by default. url-external-retrieval-args specifies the arguments that will be passed to it, ‘("-source")’ by default.

w3-netscape-compatible-comments

Not everyone uses proper html comments. To allow for the presence of lesser browsers, Emacs/W3 will honour the incorrect netscape-style comments (‘<!

>’) if w3-netscape-compatible-comments is non-nil. This is t by default,

but it shouldn’t need to be.

font-blink-interval

This controls how often blinks occur for text inside ‘<blink>’ tags. It is 0.5 seconds by default.

url-inhibit-mime-parsing

This controls whether to parse mime headers in a message. If it is nil then the headers are parsed and deleted.

url-mime-language-string

This is used to set the contents of the ‘Accept-language:’ field in http/1.0 requests. If it is nil then the field isn’t added and the server’s default language version is retrieved, if it is * then the first available langauge version is retrieved. If it is a string, then it should be the desired language.

url-multiple-p

If this is non-nil then multiple queries are possible through ‘ *URL-<i>*’ buffers.

url-personal-mail-address

url-personal-mail-address contains your full email address. This is sent in the from field in an http/1.0 request, but Chapter 7 [Security], page 49 for how to prevent this. If nil (the default), then it will be set to user-

mail-address if non-nil, else it will be (user-real-login-name) at (system-
name).
url-temporary-directory, w3-temporary-directory
url-temporary-directory and w3-temporary-directory control where tem-

porary files are placed. If ‘TMPDIR’ is set then they default to that, otherwise

‘/tmp’.
w3-documentation-root

This specifies the location of the Emacs/W3 documentation, it must end in a slash.

w3-popup-menu-on-mouse-3

If you like context-sensitive menus then you’re bound to like w3-popup-menu-on-mouse-3. If non-nil (the default) then Emacs/W3 will bind mouse-3 to provide context-sensitive menus. This might not work at the moment. If w3popup-menu-on-mouse-3 is nil, then Emacs/W3 will not change the binding of mouse-3.

w3-track-mouse

If w3-track-mouse is non-nil (the default) then Emacs/W3 will display the url under the mouse in the echo-area.

w3-use-menus

If w3-use-menus is nil then Emacs/W3 will not provide a menu interface. If it is ‘1’, then Emacs/W3 will add a ‘W3’ item to the Emacs menubar. If it is a list then Emacs/W3 will add its own menubar. The following symbols may appear in the list to control what Emacs/W3 puts in its menubar. file A list of file related commands edit Various standard editing commands (copy/paste) view Controlling various things about the document view go Navigation control bookmark Bookmark / hotlist control options Various options buffers The standard buffers menu emacs A toggle button to switch back to normal emacs menus style Control style information and who gets to set what search Various search engines help The help menu nil This may appear once in the list. All menus after this will be

displayed flush right.

12 More Help

For more help on Emacs/W3, please send me mail (wmperry+w3@cs.indiana.edu). Several discussion lists have also been created for Emacs/W3. To subscribe, send mail to majordomo@indiana.edu, with the body of the message ’subscribe listname <email addres>’. All other mail should go to <listname>@indiana.edu.

• w3-announce – this list is for anyone interested in Emacs/W3, and should in general only be used by me. The gnu.emacs.sources newsgroup and a few other mailing lists are included on this. Please only use this list for major package releases related to Emacs/W3. (www-announce@w3.org is included on this list).

• w3-beta – this list is for beta testers of Emacs/W3. These brave souls test out not-quite stable code.

• w3-dev – a list consisting of myself and a few other people who are interested in the internals of Emacs/W3, and doing active development work. Pretty dead right now, but I hope it will grow.

For help on the World Wide Web in general, please refer to the comp.infosystems.www.* newsgroups. There are also several discussion lists concerning the Web. Send mail to <listname>-request@w3.org with a subject line of ’subscribe <listname>’. All mail should go to <listname>@w3.org. Administrative mail should go to www-admin@w3.org. The lists are:

• www-talk – for general discussion of the World Wide Web, where its going, new features, etc. All the major developers are subscribed to this list.

• www-announce – for announcements concerning the World Wide Web. Server changes, new servers, new software, etc.

As a last resort, mail me. I’ll try to answer as quickly as I can.

13 Future Directions

Changes are constantly being made to the Emacs browser (hopefully all for the better). This is a list of the things that are being worked on right now.

BUGS (4.0): − need to support HTTP/0.9 (http://c2.com:8080) responses − /etc/mailcap cannot overide builtin mm-mime-data stuff? − try to protect people from using ’~’ in file URLs − keystrokes entered while in w3-pause self-insert under XEmacs — the loop around

dispatch-event needs to be smarter about what it swallows. − border-color can have multiple color specifications, but we currently choke with ’args

out of range’ when we see this. − widget appears to be stealing button3 to mean ’activate’ — this is bogus! We lose all

context-sensitive menus because of this. − We still seem to be growing the line size under Emacs 19.x/20.x − It would be really nice if w3 buffers were put into w3-mode as soon as they were created.

Then if the rendering craps out somehow then the buffer could be browsed such as it was. Ideally, links and widgets would be functional.

− document how to translate Netscape foo.pac files to emacs lisp − Should we stop using reporter.el?

BUGS (4.1): − background colors are not heeded on table rows (<tr>). Same properties on individual

cells or the table as a whole work fine. − <br> in <dd> hosed — margins in general tend to be too big sometimes. − client side imagemaps have to be in the same buffer (actually in the smae buffer,

BEFORE the usemap directive on an image) — fix to be able to use imagemaps in different files, any position, etc, etc. FEATURES (4.1)

− cache a formatted version of documents, with enough info to recreate the widgets in

them. − w3-preview-region command − LDAP support (XEmacs) − New proxy type for sending requests via mail to a mail->web->mail gateway. − Emacspeak Interaction

• some way of specifying in a stylesheet whether certain text is inaudible. use the ’inaudible text property for this.

• Full Aural-CSS support − more sophisticated filling algorithm. I’m not sure exactly what would be sufficient but

breaking lines after punctuation seems like it would solve most of the problem. − When fetching images for viewing (not inlining), W3 should at least have an option of

displaying it inline, ala Netscape.

− Widget library merging • Write a font selection widget

• Write a voice selection widget

• Write a mailcap entry widget − Custom library merging •Add custom support for MM − Hotlist handling

• Abstract out current support

• Do something similar to GNUS ’backends’ to provide easy way to add new book-

mark formats, etc. − Write a new major mode for handling CSS style sheets

FEATURES (5.0) − Emacspeak Integration

• Need option to turn off table rendering and print it out as a table that is viewable

with emacspeak-table-ui.el − Write a text/xml parser − Completely rewrite display code again

• Abstract everything out to follow parse->flow objects->render model

• Base all stylesheet stuff off of DSSSL

• CSS2

• New rendering backends − Native postscript output − LaTeX upgrade − TeXinfo

− Display code • implement <spacer> from netscape 3.0b5

• reimplement w3-show-headers

• Handle math environment using the calc library

• Better integration with the parser

Appendix A Reporting Bugs

If any bugs are discovered in Emacs/W3, please report them to the mailing list w3-beta@xemacs.org — this is where the brave souls who beta test the latest versions of Emacs/W3 reside, and are generally very responsive to bug reports. Please make sure to use the bug submission feature of Emacs/W3, so that all relevant information will be sent along with your bug report. By default this is bound to the ‘^{⟨}\underline{\overline{w}}^{⟩}’ key when in an Emacs/W3 buffer, or you can use ^{⟨}\underline{\overline{M-x w3-submit-bug}}^{⟩} from anywhere within Emacs.

For problems that are causing emacs to signal and error, please send a backtrace. You can

get a backtrace by M-x setvariable RET debug-on-error RET t RET, and then reproduce

the error.

If the problem is visual, please capture a copy of the output and mail it along with the bug report (preferably as a MIME attachment, but anything will do). You can use the xwd program under X-windows for this, or ^{⟨}\underline{\overline{Alt-PrintScreen}}^{⟩} under Windows 95/NT. Sorry, but I don’t remember what the magic incarnation is for doing a screen dump under NeXTstep or OS/2.

If the problem is actually causing Emacs to crash, then you will need to also mail the maintainers of the various Emacs distributions with the bug. Please use the gnu.emacs.bug newgroup for reporting bugs with GNU Emacs 19, and comp.emacs.xemacs for reporting bugs with XEmacs 19 or XEmacs 20. I am actively involved with the beta testing of the latest versions of both branches of Emacs, and if I can reproduce the problem, I will do my best to see it gets fixed in the next release.

It is also important to always maintain as much context as possible in your responses. I get so much email from my various Emacs-activities and work, that I cannot remember everything. If you send a bug report, and I send you a reply, and you reply with ’no that didn’t work’, then odds are I will have no clue what didn’t work, much less what that was trying to fix in the first place. It will be much quicker and less painful if I don’t have to waste a round-trip email exchange saying ’what are you talking about’.

Appendix B Dealing with Firewalls

By default, Emacs can support standard tcp/ip network connections on almost all the platforms it runs on (Unix, vms, Windows, etc). However, there are several situations where it is not sufficient.

Firewalls It is becoming more and more common to be behind a firewall or some other system that restricts your outbound network activity, especially if you are like me and away from the wonderful world of academia. Emacs/W3 has several different methods to get around firewalls (not to worry though — none of them should get you in trouble with the local mis department.)

Emacs cannot resolve hostnames.

This happens quite often on SunOS workstations and some ULTRIX machines. Some C libraries do not include the hostname resolver routines in their static libraries. If Emacs was linked statically, and was not linked with the resolver libraries, it wil not be able to get to any machines off the local network. This is characterized by being able to reach someplace with a raw ip number, but not its hostname (http://129.79.254.191/ works, but

http://www.cs.indiana.edu/ doesn’t).

The best solution for this problem is to recompile Emacs, making sure to either link dynamically (if available on your operating system), or include the

‘-lresolv’.

If you do not have the disk space or the appropriate permissions to recompile Emacs, another alternative is using the ‘nslookup’ program to do hostname resolution. To turn this on, set the variable url-gateway-broken-resolution in your ‘~/.emacs’ file. This runs the program specified by url-gateway-nslookup-program (by default "nslookup" to do hostname resolution. This program should expect a single argument on the command line — the hostname to resolve, and should produce output similar to the standard Unix ‘nslookup’ program:

     Name: www.cs.indiana.ed
     Address: 129.79.254.191

Using term (or term-like) Networking Software term^{1} for slip-like access to the internet. note: XEmacs and Emacs 19.22 or later have patches to enable native term networking. To enable it, #define TERM in the appropriate s/*.h file for the operating system, then change the SYSTEM_LIBS definition to include the ‘termnet’ library that comes with the latest versions of term. If you run into any problems with the native term networking support in Emacs or XEmacs, please let wmperry+w3@cs.indiana.edu know, as he is responsible for the original support.

1 term is a user-level protocol for emulating ip over a serial line. More information is

available at ftp://sunsite.unc.edu/pub/Linux/apps/comm/term

Emacs/W3 has support for using the gateway mechanism for certain domains, and directly connecting to others. The variable url-gateway-local-host-regexp controls this behaviour. This is a regular expression^{2} that matches local hosts that do not require the use of a gateway. If nil, then all connections are made through the gateway.

Emacs/W3 supports several methods of getting around gateways. The variable urlgateway-method controls which of these methods is used. This variable can have several values (use these as symbol names, not strings), ie: ‘(setq url-gateway-method 'telnet)’. Possible values are:

telnet Use this method if you must first telnet and log into a gateway host, and then run telnet from that host to connect to outside machines.

url-gateway-telnet-host

The gateway host to telnet to. Once logged in there, you then telnet out to the hosts you want to connect to.

url-gateway-telnet-parameters

This should be a list of parameters to pass to the ‘telnet’ program.

url-gateway-telnet-password-prompt

This is a regular expression that matches the password prompt when logging in.

url-gateway-telnet-login-prompt

This is a regular expression that matches the username prompt when logging in.

url-gateway-telnet-user-name

The username to log in with.

url-gateway-telnet-password

This is the password to send when logging in.

url-gateway-prompt-pattern

This is a regular expression that matches the shell prompt.

rlogin This method is identical to the telnet method, but uses ‘rlogin’ to log into the remote machine without having to send the username and password over the wire every time.

url-gateway-rlogin-host

Host to ‘rlogin’ to before telnetting out.

url-gateway-rlogin-parameters

Parametres to pass to ‘rsh’.

url-gateway-rlogin-user-name

User name to use when logging in to the gateway.

url-gateway-prompt-pattern

This is a regular expression that matches the shell prompt.

2 Please see the full Emacs distribution for a description of regular expressions tcp Masanobu UMEDA (umerin@mse.kyutech.ac.jp) has written a very small ap-

plication that you can run in a subprocess to do the network connections.

socks Use if the firewall has a socks gateway running on it. socks v5 protocol is defined in rfc1928.

socks-password

If this is nil then you will be asked for the passward, otherwise it will be used as the password for authenticating you to the socks server.

socks-username

This is the username to use when authenticating yourself to the socks server. By default this is your login name

socks-timeout

This controls how long, in seconds, Emacs/W3 will wait for responses from the socks server; it is 5 by default.

socks-server

Thiss the default server, it take the form (‘"Default server"’ server port version) where version can be either 4 or 5.

socks-server-aliases

This a list of server aliases. It is a list of aliases of the form (alias hostname port version).

socks-network-aliases

This a list of network aliases. Each entry in the list takes the form (alias (network)) where alias is a string that names the network. The networks can contain a pair (not a dotted pair) of ip addresses which specify a range of ip addresses, an ip address and a netmask, a domain name or a unique hostname or ip address.

socks-redirection-rules

This a list of redirection rules. Each rule take the form (Destination network Connection type) where Destination network is a network alias from socks-network-aliases and Connection type can be nil in which case a direct connection is used, or it can be an alias from socks-server-aliases in which case that server is used as a proxy.

socks-nslookup-program

This the ‘nslookup’ program. It is ‘nslookup’ by default.

native This means that Emacs/W3 should use the builtin networking code of Emacs. This should be used only if there is no firewall, or the Emacs source has already been hacked to get around the firewall.

Emacs/W3 should now be able to get outside the local network. If none of this makes sense, its probably my fault. Please check with the network administrators to see if they have a program that does most of this already, since somebody somewhere at the company has probably been through something similar to this before, and would be much more Appendix B: Dealing with Firewalls 66 helpful/knowledgeable about the local setup than I would be. But feel free to mail me as a last resort.

Appendix C Proxy Gateways

In late January 1993, Kevin Altis and Lou Montulli proposed and implemented a new proxy service. This service requires the use of environment variables to specify a gateway server/port # to send protocol requests to. Each protocol (http, wais, gopher, ftp, etc.) can have a different gateway server. The environment variables are PROTOCOL proxy, where PROTOCOL is one of the supported network protocols (gopher, file, http, ftp, etc.)

For companies with internal intranets, it will usually be helpful to define a list of hosts that should be contacted directly, not sent through the proxy. The NO_PROXY environment variable controls what hosts are able to be contacted directly. This should be a comma separated list of hostnames, domain names, or a mixture of both. Asterisks can be used as a wildcard. For example:

     NO_PROXY=*.aventail.com,home.com,*.seanet.com

tells Emacs/W3 to contact all machines in the aventail.com and seanet.com domains directly, as well as the machine named home.com.

For those adventurous souls who enjoy writing regular expressions, all the proxy settings can be manipulated from Emacs-Lisp. The variable url-proxy-services controls this. This is an assoc list, keyed on the protocol type (http, gopher, etc) in all lowercase. The cdr of each entry should be the address of the proxy server to contact, followed by ":" and the port number to use. In the case of the special "no proxy" entry, it should be a regular expression that matches any hostnames that should be contacted directly.

     (setq url-proxy-services
            '(("http"     . "proxy.aventail.com:80")
              ("no_proxy" . "^.*\\(aventail\\|seanet\\)\.com")))

Appendix D Installing SSL

In order to use SSL in Emacs/W3, an implementation of SSL is necessary. Emacs/W3 is configued to work out of the box with SSLeay 0.6.6 or later. For best results, you should apply a patch that makes the SSLeay client much quieter about what it reports.

   You can download SSLeay from ftp://ftp.psy.uq.oz.au/pub/Crypto/SSL/

The following variables control how the external program is invoked.

ssl-program-name

The name of the program to run, as a string.

               (setq ssl-program-name "s_client")
ssl-program-arguments

This should be used if your SSL program needs command line switches to specify any behaviour (certificate file locations, etc). This is a list of strings and symbols. The special symbols ’host and ’port may be used in the list of arguments and will be replaced with the hostname and service/port that will be connected to.

               (setq ssl-program-arguments '("-host" host
                                             "-port" service
                                             "-verify" "4"
                                             "-CApath /usr/local/ssl/certs"))

The default is ("-host" host "-port" service "-verify" ssl-certificate-verification-policy -CApath ssl-certificate-directory).

ssl-certificate-directory is the directory in which ca certificates are stored. It is w3-configuration-directory/cert by default.

ssl-rehash-program-name is the program that is run after adding a certificate to the ssl-certificate-directory directory. It is run with the directory name as an argument and defaults to c_rehash.

ssl-view-certificate-program-name names the program that can produce a humanreadable view of a certificate. It is x509 by default and is called with the arguments listed

in ssl-view-certificate-program-arguments which is ("text" "-inform" "DER") by

default.

ssl-certificate-directory-style specifies the type of certificate database to use. It’s default (and at the moment, only possible value) is ssleay which specifies a directory or pem encoded certificates with hash symlinks.

You can decide how high up the chain of certificates should be verified by setting ssl-

certificate-verification-policy. Possible values are

0 No verification

1 Verification required

3 Reject connection if verification fails

5 SSL VERIFY CLIENT ONCE

The default is 0

Appendix E Mailcap Files

NCSA Mosaic and almost all other WWW browsers rely on a separate file for mapping MIME types to external viewing programs. This takes some of the burden off of browser developers, so each browser does not have to support all image formats, or postscript, etc. Instead of having the users of Emacs/W3 duplicate this in lisp, this file can be parsed using the mm-parse-mailcaps function. This function is called each time Emacs/W3 is loaded. It tries to locate mimetype files in several places. If the environment variable MAILCAPS is nonempty, then this is assumed to specify a UNIX-like path of mimetype files (this is a colon separated string of pathnames). If the MAILCAPS environment variable is empty, then Emacs/W3 looks for these files:

 1. ‘~/.mailcap’
 2. ‘/etc/mailcap’
 3. ‘/usr/etc/mailcap’
 4. ‘/usr/local/etc/mailcap’

This format of this file is specified in RFC 1343, but a brief synopsis follows (this is taken verbatim from sections of RFC 1343).

Each mailcap file consists of a set of entries that describe the proper handling of one media type at the local site. For example, one line might tell how to display a message in Group III fax format. A mailcap file consists of a sequence of such individual entries, separated by newlines (according to the operating system’s newline conventions). Blank lines and lines that start with the "#" character (ASCII 35) are considered comments, and are ignored. Long entries may be continued on multiple lines if each non-terminal line ends with a backslash character (’\’, ASCII 92), in which case the multiple lines are to be treated as a single mailcap entry. Note that for such "continued" lines, the backslash must be the last character on the line to be continued.

Each mailcap entry consists of a number of fields, separated by semi-colons. The first two fields are required, and must occur in the specified order. The remaining fields are optional, and may appear in any order.

The first field is the content-type, which indicates the type of data this mailcap entry describes how to handle. It is to be matched against the type/subtype specification in the "Content-Type" header field of an Internet mail message. If the subtype is specified as "*", it is intended to match all subtypes of the named content-type.

The second field, view-command, is a specification of how the message or body part can be viewed at the local site. Although the syntax of this field is fully specified, the semantics of program execution are necessarily somewhat operating system dependent.

The optional fields, which may be given in any order, are as follows:

• The "compose" field may be used to specify a program that can be used to compose a new body or body part in the given format. Its intended use is to support mail composing agents that support the composition of multiple types of mail using external composing agents. As with the view- command, the semantics of program execution are operating system dependent. The result of the composing program may be data that is not yet suitable for mail transport—that is, a Content-Transfer-Encoding may need to be applied to the data.

• The "composetyped" field is similar to the "compose" field, but is to be used when the composing program needs to specify the Content-type header field to be applied to the composed data. The "compose" field is simpler, and is preferred for use with existing (non-mail-oriented) programs for composing data in a given format. The "composetyped" field is necessary when the Content-type information must include auxilliary parameters, and the composition program must then know enough about mail formats to produce output that includes the mail type information.

• The "edit" field may be used to specify a program that can be used to edit a body or body part in the given format. In many cases, it may be identical in content to the "compose" field, and shares the operating-system dependent semantics for program execution.

• The "print" field may be used to specify a program that can be used to print a message or body part in the given format. As with the view-command, the semantics of program execution are operating system dependent.

• The "test" field may be used to test some external condition (e.g. the machine architecture, or the window system in use) to determine whether or not the mailcap line applies. It specifies a program to be run to test some condition. The semantics of execution and of the value returned by the test program are operating system dependent. If the test fails, a subsequent mailcap entry should be sought. Multiple test fields are not permitted—since a test can call a program, it can already be arbitrarily complex.

• The "needsterminal" field indicates that the view-command must be run on an interactive terminal. This is needed to inform window-oriented user agents that an interactive terminal is needed. (The decision is not left exclusively to the view-command because in some circumstances it may not be possible for such programs to tell whether or not they are on interactive terminals.) The needsterminal command should be assumed to apply to the compose and edit commands, too, if they exist. Note that this is NOT a test—it is a requirement for the environment in which the program will be executed, and should typically cause the creation of a terminal window when not executed on either a real terminal or a terminal window.

• The "copiousoutput" field indicates that the output from the view-command will be an extended stream of output, and is to be interpreted as advice to the UA (User Agent mail- reading program) that the output should be either paged or made scrollable. Note that it is probably a mistake if needsterminal and copiousoutput are both specified.

• The "description" field simply provides a textual description, optionally quoted, that describes the type of data, to be used optionally by mail readers that wish to describe the data before offering to display it.

• The "x11-bitmap" field names a file, in X11 bitmap (xbm) format, which points to an appropriate icon to be used to visually denote the presence of this kind of data.

• Any other fields beginning with "x-" may be included for local or mailer-specific extensions of this format. Implementations should simply ignore all such unrecognized fields to permit such extensions, some of which might be standardized in a future version of this document.

Appendix F: General Index 71

Appendix F General Index

.

.w3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3

32-Bit Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

A

Authentication, Basic . . . . . . . . . . . . . . . . . . . . . . . . 49 Authentication, Digest . . . . . . . . . . . . . . . . . . . . . . . 49 AXP-VMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

B

Bad HTML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Basic Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Broken SunOS libc. . . . . . . . . . . . . . . . . . . . . . . . . . . 63 browse-url-browser-function . . . . . . . . . . . . . . . . . . 54 Browser emulation . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Browser masquerading . . . . . . . . . . . . . . . . . . . . . . . 13 Browsing with no network connection . . . . . . . . . 53 Buffer movement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Bugs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

C

Cache cleaning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Cache only mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Caching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Cleaning the cache . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Clearing the cache . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Clueless in Seattle . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Completion of URLs . . . . . . . . . . . . . . . . . . . . . . . . . 15 Contacting the author. . . . . . . . . . . . . . . . . . . . . . . . 62 Cookies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

D

data urls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Default stylesheet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Digital VMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Disk Cache . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 DOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

E

Emulation of other browsers . . . . . . . . . . . . . . . . . . 10 Export Restrictions. . . . . . . . . . . . . . . . . . . . . . . 49, 68 Exportability . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 68

F

Faulty hostname resolvers . . . . . . . . . . . . . . . . . . . . 63 File urls. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 finger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 finger urls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Firewalls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 font-blink-interval. . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 ftp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

G

Gag Puke Retch. . . . . . . . . . . . . . . . . . . . . . . . . . 49, 68 Getting Started. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 GNUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 gopher . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 gopher urls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

H

history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 History Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Hostname resolution . . . . . . . . . . . . . . . . . . . . . . . . . 63 http . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 HTTP Proxy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 HTTP/1.0 Authentication . . . . . . . . . . . . . . . . 49, 68 https . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

I

info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 info urls. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Interfacing to Mail/News . . . . . . . . . . . . . . . . . . . . . 54 Invalid HTML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 irc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 irc urls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

L

LaTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Limiting the size of the cache . . . . . . . . . . . . . . . . . 53 Lynx emulation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

M

mailserver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 mailserver urls. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 mailto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 mailto urls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Masquerading as other browsers . . . . . . . . . . . . . . 13 Microsloth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 mime-types file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 mm-add-mailcap-entry. . . . . . . . . . . . . . . . . . . . . . . 48 mm-content-transfer-encodings. . . . . . . . . . . . . . . . 56 mm-mime-extensions . . . . . . . . . . . . . . . . . . . . . . . . . 47 mm-parse-mimetypes . . . . . . . . . . . . . . . . . . . . . . . . . 47 MS-DOG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 MS-DOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

N

Netless browsing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Netscape emulation . . . . . . . . . . . . . . . . . . . . . . . . . . 12 news . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 news urls. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Newsgroups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 nfs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 nfs urls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 nntp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 No Proxy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 NO PROXY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Non-Unix Operating Systems . . . . . . . . . . . . . . . . . 51

O

OS/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

P

Paranoia . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Persistent Cache . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Postscript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Printing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 profile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Proxies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Proxies, environment variables . . . . . . . . . . . . . . . . 67 Proxies, exclusion lists . . . . . . . . . . . . . . . . . . . . . . . 67 Proxies, setting from lisp . . . . . . . . . . . . . . . . . . . . . 67 Proxy Servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

R

Relevant Newsgroups . . . . . . . . . . . . . . . . . . . . . . . . 59 Relying on cache. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Reporting Bugs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 rlogin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 rlogin urls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 RMAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

S

scroll-down . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Secure Sockets Layer . . . . . . . . . . . . . . . . . . . . . 49, 68 Security. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Security, Basic. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Security, Digest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 snews . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 socks-network-aliases . . . . . . . . . . . . . . . . . . . . . . . . . 65 socks-nslookup-program . . . . . . . . . . . . . . . . . . . . . . 65 socks-password . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 socks-redirection-rules . . . . . . . . . . . . . . . . . . . . . . . . 65 socks-server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 socks-server-aliases. . . . . . . . . . . . . . . . . . . . . . . . . . . 65 socks-timeout. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 socks-username . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 SSL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 68 ssl-certificate-directory . . . . . . . . . . . . . . . . . . . . . . . 68 ssl-certificate-directory-style . . . . . . . . . . . . . . . . . . 68 ssl-certificate-verification-policy . . . . . . . . . . . . . . . 68 ssl-program-arguments . . . . . . . . . . . . . . . . . . . . . . . 68 ssl-program-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 ssl-rehash-program-name . . . . . . . . . . . . . . . . . . . . . 68 ssl-view-certificate-program-arguments. . . . . . . . . 68 ssl-view-certificate-program-name . . . . . . . . . . . . . 68 Standalone mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Startup files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

T

telnet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 telnet urls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 term . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 tn3270 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 tn3270 urls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

turn-off-arena-masquerade-mode . . . . . . . . . . . . 13
turn-off-ie-masquerade-mode . . . . . . . . . . . . . . . 13
turn-off-lynx-masquerade-mode . . . . . . . . . . . . . 13
turn-off-netscape-masquerade-mode. . . . . . . . . 13
turn-on-arena-masquerade-mode . . . . . . . . . . . . . 13
turn-on-ie-masquerade-mode . . . . . . . . . . . . . . . . 13
turn-on-lynx-emulation . . . . . . . . . . . . . . . . . . . . 10
turn-on-lynx-masquerade-mode . . . . . . . . . . . . . . 13
turn-on-netscape-emulation . . . . . . . . . . . . . . . . 12
turn-on-netscape-masquerade-mode. . . . . . . . . . 13

Turning on caching . . . . . . . . . . . . . . . . . . . . . . . . . . 53

U

url-automatic-caching . . . . . . . . . . . . . . . . . . . . . . . . 53 url-bad-port-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 url-be-anal-about-file-attributes . . . . . . . . . . . . . . . 44 url-be-asynchronous. . . . . . . . . . . . . . . . . . . . . . . . . . 17 url-cache-creation-function. . . . . . . . . . . . . . . . . . . . 53 url-cache-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 url-cache-expired . . . . . . . . . . . . . . . . . . . . . . . . . . 53 url-cache-ignored-protocols . . . . . . . . . . . . . . . . . . . 53 url-confirmation-func. . . . . . . . . . . . . . . . . . . . . . . . . 56 url-cookie-confirmation . . . . . . . . . . . . . . . . . . . . . . . 50 url-cookie-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 url-cookie-trusted-urls . . . . . . . . . . . . . . . . . . . . . . . . 50 url-cookie-untrusted-urls . . . . . . . . . . . . . . . . . . . . . 50 url-default-ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 url-default-retrieval-proc . . . . . . . . . . . . . . . . . . . . . 17 url-directory-index-file . . . . . . . . . . . . . . . . . . . . . . . 43 url-external-retrieval-args. . . . . . . . . . . . . . . . . . . . . 57 url-external-retrieval-program. . . . . . . . . . . . . . . . . 57 url-gateway-broken-resolution . . . . . . . . . . . . . . . . . 63 url-gateway-local-host-regexp . . . . . . . . . . . . . . . . . 64 url-gateway-method . . . . . . . . . . . . . . . . . . . . . . . . . . 64 url-gateway-prompt-pattern . . . . . . . . . . . . . . . . . . 64 url-gateway-rlogin-host . . . . . . . . . . . . . . . . . . . . . . . 64 url-gateway-rlogin-parameters . . . . . . . . . . . . . . . . 64 url-gateway-rlogin-user-name . . . . . . . . . . . . . . . . . 64 url-gateway-telnet-host . . . . . . . . . . . . . . . . . . . . . . . 64 url-gateway-telnet-login-prompt . . . . . . . . . . . . . . . 64 url-gateway-telnet-parameters . . . . . . . . . . . . . . . . 64 url-gateway-telnet-password . . . . . . . . . . . . . . . . . . 64 url-gateway-telnet-password-prompt . . . . . . . . . . . 64 url-gateway-telnet-user-name . . . . . . . . . . . . . . . . . 64 url-global-history-file . . . . . . . . . . . . . . . . . . . . . . 4, 14 url-global-history-save-interval . . . . . . . . . . . . . . 4, 14 url-gopher-icons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 url-gopher-labels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 url-gopher-to-mime . . . . . . . . . . . . . . . . . . . . . . . . . . 45 url-honor-refresh-requests. . . . . . . . . . . . . . . . . . . . . 44 url-inhibit-mime-parsing. . . . . . . . . . . . . . . . . . . . . . 57 url-irc-function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 url-keep-history . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 14 url-list-processes . . . . . . . . . . . . . . . . . . . . . . . . . . 8 url-mail-command . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 url-max-password-attempts . . . . . . . . . . . . . . . . . . . 56 url-mime-language-string . . . . . . . . . . . . . . . . . . . . . 57 url-multiple-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 url-news-server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 url-news-use-article-mode. . . . . . . . . . . . . . . . . . . . . 45 url-nfs-automounter-directory-spec . . . . . . . . . . . . 43 url-passwd-entry-func . . . . . . . . . . . . . . . . . . . . . . . . 56 url-personal-mail-address . . . . . . . . . . . . . . . . . . . . . 57 url-privacy-level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 url-proxy-services . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 url-show-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 url-standalone-mode . . . . . . . . . . . . . . . . . . . . . . . . . 53 url-temporary-directory . . . . . . . . . . . . . . . . . . . . . . 57 url-uncompressor-alist . . . . . . . . . . . . . . . . . . . . . . . . 56 url-use-hypertext-dired . . . . . . . . . . . . . . . . . . . . . . . . 7 url-use-hypertext-gopher . . . . . . . . . . . . . . . . . . . . . 45 url-view-url . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Usefulness of global history . . . . . . . . . . . . . . . . . . . 15 Using Emacs/W3 with Gnus. . . . . . . . . . . . . . . . . . 54 Using Emacs/W3 with RMAIL . . . . . . . . . . . . . . . 54 Using Emacs/W3 with VM . . . . . . . . . . . . . . . . . . . 54

V

VAX-VMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 VM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 VMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

W

w3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

w3-arena-masquerade-mode . . . . . . . . . . . . . . . . . . 13

w3-auto-image-alt . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

w3-backward-in-history . . . . . . . . . . . . . . . . . . 5, 14

w3-bullets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 w3-complete-link. . . . . . . . . . . . . . . . . . . . . . . . . . 6, 7 w3-configuration-directory . . . . . . . . . . . . . . . . . . . . . 4 w3-debug-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 w3-debug-html . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 w3-default-configuration-file . . . . . . . . . . . . . . . . . . . 4 w3-default-homepage. . . . . . . . . . . . . . . . . . . . . . . . . . 2

w3-default-image-alt-func . . . . . . . . . . . . . . . . . 16

w3-default-stylesheet . . . . . . . . . . . . . . . . . . . . . . . . . . 4 w3-defined-link-types. . . . . . . . . . . . . . . . . . . . . . . . . 16 w3-delay-image-loads. . . . . . . . . . . . . . . . . . . . . . . . . 16 w3-display-errors-hook . . . . . . . . . . . . . . . . . . . . . . . 55 w3-display-frames. . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 w3-do-incremental-display . . . . . . . . . . . . . . . . . . . . 17 w3-do-scripting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

w3-document-information . . . . . . . . . . . . . . . . . . . . 6

w3-documentation-root. . . . . . . . . . . . . . . . . . . . . . . 57 w3-download-this-url. . . . . . . . . . . . . . . . . . . . . . . . 7

w3-download-url-at-point . . . . . . . . . . . . . . . . . . . 7

w3-dummy-img-alt-repl . . . . . . . . . . . . . . . . . . . . . . 16 w3-echo-link. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 w3-end-of-document . . . . . . . . . . . . . . . . . . . . . . . . . . 5 w3-fetch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 14

w3-follow-inlined-image . . . . . . . . . . . . . . . . . . . . 7

w3-follow-link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 w3-follow-mouse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 w3-forward-in-history . . . . . . . . . . . . . . . . . . . 5, 14 w3-goto-last-buffer . . . . . . . . . . . . . . . . . . . . . . 6, 7 w3-help. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 w3-history-backward . . . . . . . . . . . . . . . . . . . . . . . . . 5 w3-history-forward . . . . . . . . . . . . . . . . . . . . . . . . . . 5 w3-honor-stylesheets . . . . . . . . . . . . . . . . . . . . . . . . . 17 w3-horizontal-rule-char . . . . . . . . . . . . . . . . . . . . . . . 15

w3-hotlist-add-document . . . . . . . . . . . . . . . . . . . 13
w3-hotlist-add-document-at-point. . . . . . . . . . 14

w3-hotlist-append . . . . . . . . . . . . . . . . . . . . . . . . . . 14 w3-hotlist-apropos . . . . . . . . . . . . . . . . . . . . . . . . . 13 w3-hotlist-delete . . . . . . . . . . . . . . . . . . . . . . . . . . 13 w3-hotlist-file . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 13, 14 w3-hotlist-refresh . . . . . . . . . . . . . . . . . . . . . . . . . 14

w3-hotlist-rename-entry . . . . . . . . . . . . . . . . . . . 13

w3-html-errors-font-lock-keywords. . . . . . . . . . . . . 55 w3-icon-format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 w3-ie-masquerade-mode . . . . . . . . . . . . . . . . . . . . . 13 w3-image-mappings . . . . . . . . . . . . . . . . . . . . . . . . . . 17

w3-insert-formatted-url . . . . . . . . . . . . . . . . . . . . 8

w3-insert-this-url . . . . . . . . . . . . . . . . . . . . . . . . . . 8 w3-latex-docstyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 w3-latex-packages. . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 w3-latex-print-links . . . . . . . . . . . . . . . . . . . . . . . . . . 54 w3-latex-use-latex2e . . . . . . . . . . . . . . . . . . . . . . . . . 54 w3-latex-use-maketitle . . . . . . . . . . . . . . . . . . . . . . . 54 w3-leave-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 8 w3-load-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

w3-lynx-emulation-minor-mode . . . . . . . . . . . . . . 10
w3-lynx-masquerade-mode . . . . . . . . . . . . . . . . . . . 13
w3-mail-current-document . . . . . . . . . . . . . . . . . . . 8
w3-mail-document-author . . . . . . . . . . . . . . . . . . . . 8
w3-mail-document-under-point . . . . . . . . . . . . . . . 8

w3-max-menu-length . . . . . . . . . . . . . . . . . . . . . . . . . 17 w3-max-menu-width . . . . . . . . . . . . . . . . . . . . . . . . . 17 w3-maximum-line-length . . . . . . . . . . . . . . . . . . . . . 17 w3-min-img-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 w3-mode-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 w3-modeline-format . . . . . . . . . . . . . . . . . . . . . . . . . . 17 w3-netscape-compatible-comments . . . . . . . . . . . . 57

w3-netscape-emulation-minor-mode. . . . . . . . . . 12
w3-netscape-masquerade-mode . . . . . . . . . . . . . . . 13

w3-next-document . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 w3-notify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 w3-open-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 w3-popup-info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 w3-popup-menu-on-mouse-3 . . . . . . . . . . . . . . . . . . 57 w3-postscript-print-function . . . . . . . . . . . . . . . . . . 54 w3-prev-document . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 w3-print-command . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 w3-print-this-url . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

w3-print-url-under-point . . . . . . . . . . . . . . . . . . . 7

w3-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 8 w3-refresh-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 w3-reload-document . . . . . . . . . . . . . . . . . . . . . . . . . . 7 w3-reuse-buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 w3-right-margin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 w3-save-as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 w3-save-this-url . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 w3-save-url . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 w3-scroll-up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 w3-show-history-list . . . . . . . . . . . . . . . . . . . . 7, 14 w3-show-hotlist . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 w3-source-document . . . . . . . . . . . . . . . . . . . . . . . . . . 6

w3-source-document-at-point . . . . . . . . . . . . . . . . 6

w3-start-of-document. . . . . . . . . . . . . . . . . . . . . . . . 5 w3-submit-bug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 w3-temporary-directory . . . . . . . . . . . . . . . . . . . . . . 57 w3-track-mouse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 w3-use-hotlist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 w3-use-links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 w3-use-menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 w3-use-terminal-characters . . . . . . . . . . . . . . . . . . . 16 w3-use-terminal-characters-on-tty . . . . . . . . . . . . . 16 w3-use-terminal-glyphs . . . . . . . . . . . . . . . . . . . . . . . 16 w3-user-colors-take-precedence . . . . . . . . . . . . . . . . 17 w3-user-fonts-take-precedence. . . . . . . . . . . . . . . . . 17 w3-version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 w3-view-this-url . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 w3-widget-backward . . . . . . . . . . . . . . . . . . . . . . . . . . 5 w3-widget-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Warp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Windows ’95 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Windows (32-Bit). . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

Appendix G: Key Index 76

Appendix G Key Index

!_{! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11} #_{# . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10} (_{( . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10} )_{) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10} *_{* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11} -_{- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10} /_{/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11} =_{= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11} ?_{? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 11} [_{[ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11} ]_{] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11} +_{+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10} >

> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

\

\ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

<

< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

A

a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 A-t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

B

b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 10 B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 backspace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 11

C

c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 11 C-? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 C-a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-c C-b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 C-k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 C-n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 C-t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 C-u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 C-w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

Control Middle Mouse Button . . . . . . . . . . . . . . . . . 7

D

d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 11 D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 down . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 12

E

e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

F

f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

G

g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 11

H

h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 H F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ha . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 hA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 HB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 hd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 hi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 hI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 hr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 hR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 hu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 hv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Hv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

I

i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 11 I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

J

j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

K

k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 11 K . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

L

l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 7, 11 left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 12

M

m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 7, 11 M-a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-del . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 M-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-m. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 12 M-M . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 M-n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-right. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-s. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 12 M-space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 M-t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 M-tab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 8 M-x w3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

M-x w3-insert-formatted-url . . . . . . . . . . . . . . . . 8

Meta return . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Meta-tab. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Middle Mouse Button . . . . . . . . . . . . . . . . . . . . . . . . . 6

N

n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 11

O

o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 11

P

p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 11 P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

Q

q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 8, 11 Q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Q, u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

R

r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 11 R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 return . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 6, 10 right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 12

S

s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 11 S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 10

T

tab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

U

u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 11 U . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 up. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 12

V

v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 11 V . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 11

W

w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

X

x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

Z

z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

Table of Contents

W3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Getting Started. . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

1.1 Downloading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Building and Installing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Startup Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

1.3.1 Emacs/W3 profile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.2 Default stylesheets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.3 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.4 Hotlists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Basic Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2.1 Movement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2 Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3 Action. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.4 Miscellaneous . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

3 Compatibility with other Browsers. . . . . . . . . 10

3.1 Emulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.2 Lynx emulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.3 Netscape emulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.4 Masquerading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.5 Hotlist Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.6 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.7 Global History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.8 Display Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

3.8.1 Asynchronous behaviour . . . . . . . . . . . . . . . . . . . . . . . 17

4 Stylesheets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

4.1 Terminology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 4.2 Basic Concepts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.3 Pseudo-Classes/Elements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.4 The Cascade . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.5 Properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

4.5.1 Font Properties. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.5.1.1 font-family . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.5.1.2 font-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.5.1.3 font-variant. . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.5.1.4 font-weight . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 4.5.1.5 font-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 4.5.1.6 font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

4.5.2 Colors and Backgrounds. . . . . . . . . . . . . . . . . . . . . . . . 29 4.5.2.1 color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.5.2.2 background-color. . . . . . . . . . . . . . . . . . . . . . 30 4.5.2.3 background-image. . . . . . . . . . . . . . . . . . . . . 30 4.5.2.4 background-repeat . . . . . . . . . . . . . . . . . . . . 30 4.5.2.5 background-attachment. . . . . . . . . . . . . . . . 30 4.5.2.6 background-position . . . . . . . . . . . . . . . . . . . 30 4.5.2.7 background . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

4.5.3 Text Properties. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 4.5.3.1 word-spacing . . . . . . . . . . . . . . . . . . . . . . . . . 31 4.5.3.2 letter-spacing . . . . . . . . . . . . . . . . . . . . . . . . . 31 4.5.3.3 text-decoration . . . . . . . . . . . . . . . . . . . . . . . 32 4.5.3.4 vertical-align . . . . . . . . . . . . . . . . . . . . . . . . . 32 4.5.3.5 text-transform . . . . . . . . . . . . . . . . . . . . . . . . 32 4.5.3.6 text-align. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.5.3.7 text-indent . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.5.3.8 line-height . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

4.5.4 Box Properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.5.5 Classification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

4.5.5.1 display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.5.5.2 white-space . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 4.5.5.3 list-style-type . . . . . . . . . . . . . . . . . . . . . . . . . 34 4.5.5.4 list-style-image. . . . . . . . . . . . . . . . . . . . . . . . 35 4.5.5.5 list-style-position. . . . . . . . . . . . . . . . . . . . . . 35 4.5.5.6 list-style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

4.5.6 Media Selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.5.7 Speech Properties. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

4.5.7.1 volume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.5.7.2 pause-before . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.5.7.3 pause-after . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.5.7.4 pause . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.5.7.5 cue-before . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.5.7.6 cue-after . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.5.7.7 cue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.5.7.8 cue-during . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.5.7.9 speed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.5.7.10 voice-family . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.5.7.11 pitch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.5.7.12 pitch-range . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.5.7.13 stress . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.5.7.14 richness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.5.7.15 speak-punctuation . . . . . . . . . . . . . . . . . . . 41 4.5.7.16 speak-date. . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.5.7.17 speak-numeral . . . . . . . . . . . . . . . . . . . . . . . 41 4.5.7.18 speak-time . . . . . . . . . . . . . . . . . . . . . . . . . . 41

4.6 Units. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.6.1 Length Units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.6.2 Percentage Units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.6.3 color Units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.6.4 URLs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.6.5 Angle Units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.6.6 Time Units. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

5 Supported URLs. . . . . . . . . . . . . . . . . . . . . . . . . . 43

5.1 File urls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.2 ftp urls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.3 nfs urls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.4 info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.5 http/https . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.6 mailto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.7 mailserver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.8 news/nntp/snews . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.9 rlogin/telnet/tn3270 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.10 irc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.11 data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.12 gopher . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.13 finger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 5.14 netrek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

6 MIME Support . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

6.1 Adding MIME types based on file extensions . . . . . . . . . . . . . 47 6.2 Specifying Viewers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

7 Security . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

7.1 Privacy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

8 Cookies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

9 Non-Unix Operating Systems. . . . . . . . . . . . . . 51

9.1 VMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 9.2 OS/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 9.3 MS-DOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 9.4 Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

10 Speech Integration. . . . . . . . . . . . . . . . . . . . . . . 52

11 Advanced Features . . . . . . . . . . . . . . . . . . . . . . 53

11.1 Disk Caching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 11.2 Printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 11.3 Interfacing to Mail/News . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 11.4 Debugging HTML. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 11.5 Hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 11.6 Miscellaneous variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

12 More Help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 13 Future Directions . . . . . . . . . . . . . . . . . . . . . . . . 60 Appendix A Reporting Bugs . . . . . . . . . . . . . . . . 62 Appendix B Dealing with Firewalls . . . . . . . . . . 63 Appendix C Proxy Gateways. . . . . . . . . . . . . . . . 67 Appendix D Installing SSL . . . . . . . . . . . . . . . . . . 68 Appendix E Mailcap Files. . . . . . . . . . . . . . . . . . . 69 Appendix F General Index . . . . . . . . . . . . . . . . . . 71 Appendix G Key Index . . . . . . . . . . . . . . . . . . . . . 76